### v1

# FuncGenFoil: Airfoil Generation and Editing Model in Function Space

Abstract
Aircraft manufacturing is the jewel in the crown of industry, among which generating high-fidelity airfoil geometries with controllable and editable representations remains a fundamental challenge. While existing deep-learning-based methods rely on predefined parametric function families, e.g., Bézier curves and discrete point-based representations, they suffer from inherent trade-offs between expressiveness and resolution flexibility. To tackle this challenge, we introduce FuncGenFoil, a novel function-space generative model that directly learns functional airfoil geometries. Our method inherits both the advantages of arbitrary resolution sampling and the smoothness of parametric functions, as well as the strong expressiveness of discrete point-based functions. Empirical evaluations on the AFBench dataset demonstrate that FuncGenFoil improves upon state-of-the-art methods in airfoil generation by achieving a relative -74.4% label error reduction and +23.2% diversity increase on the AF-200K dataset. Our results highlight the advantages of function-space modeling for aerodynamic shape optimization, offering a powerful and flexible framework for high-fidelity airfoil design. Our code will be released.
1 Introduction
Figure 1: The conceptual difference between FuncGenFoil and previous airfoil representation methods. In many previous approaches, airfoils are represented either as parametric models, as shown in (a), or as discrete point-based models, as shown in (b). In contrast, (c) illustrates FuncGenFoil’s approach, where an airfoil is treated as a continuous function mapped from a latent function, enabling a generative model in function space.
The airfoil inverse design problem serves as a central aspect of aircraft manufacturing. Traditionally, given geometric requirements, engineers first select the most similar airfoils from well-known airfoil datasets (e.g., NACA Cummings et al. (2015)) and leverage the trial-and-error strategy Sharma et al. (2021). Considering the mission of the aircraft, an initial airfoil design that meets the design conditions is preliminarily created. Then, through rounds of physical analysis, such as aerodynamics and mechanics, the airfoil is iteratively optimized to achieve better performance until the requirements are met. In practice, such direct design procedures are highly inefficient and time-consuming, often taking months. To minimize development and design time, as well as associated costs, automatic design methods have been introduced as efficient alternatives in aircraft manufacturing engineering. In particular, machine learning-based design and optimization techniques have gained significant attention. However, before applying algorithms to airfoil design, it is crucial to determine appropriate methods for representing airfoils within these algorithms.
Existing methods for airfoil representation can generally be divided into two categories: parametric-model-based approaches Xie et al. (2024) and discrete-point-based methods Liu et al. (2024); Sekar et al. (2019). First, parametric-model-based methods predefine function families, e.g., Bézier curves Chen and Fuge (2021), Hicks-Henne curves, and NURBS, and leverage mathematical optimization techniques or generative models to determine the parameters of these functions for a new airfoil. These methods rigorously preserve key geometric properties of the defined functional families, e.g., high-order smoothness. Furthermore, such functional representations of airfoils allow for arbitrary sampling of control points in real manufacturing, given the constraints of engineering precision. Despite these benefits, parametric-model-based methods suffer from a significantly reduced design space, i.e., selecting a specific function family excludes the possibility of other shapes, which limits the upper bound of airfoil design algorithms. Second, discrete-point-based methods directly generate multiple points to represent airfoil shapes. These methods maximize the design space of airfoils but cannot maintain some important mathematical properties, e.g., continuity. Furthermore, they cannot directly generate control points at arbitrary resolutions because the number of generated points is typically fixed for each model after training.
To address the trade-offs between these two mainstream approaches, we ask: Can we design an algorithm that leverages the advantages of both?
In this paper, we answer this question by proposing FuncGenFoil, a novel function-space generative model for airfoil representation (see Fig. 1). Different from previous data-driven and generative methods, e.g., cVAE Kingma (2013) and cGAN Mirza (2014), which directly generate discrete points, our approach models airfoil geometry using a general continuous function approximator Anandkumar et al. (2019); Li et al. (2021); Kovachki et al. (2023); Azizzadenesheli et al. (2024), i.e., neural operator architectures. Simultaneously, it leverages recent advances in generative methods, e.g., diffusion models Ho et al. (2020a) and flow matching models Lipman et al. (2022), to generate diverse airfoils beyond the design space of any predefined geometry function. Our method has the advantages of both parametric-model-based approaches and discrete-point-based methods. Due to its functional representation, the generated airfoil is continuous and can be sampled at arbitrary resolutions, making it easier to manufacture. Moreover, thanks to the general functional approximator nature of neural operators, our method can explore a broader design space beyond predefined functional families.
Specifically, we leverage the flow matching framework Lipman et al. (2022), an improved alternative to diffusion models, and FNO Li et al. (2021), a resolution-free neural operator, as the backbone of our generative model to design FuncGenFoil. In the forward process, we perturb the airfoil into a noise distribution through straight flows. In the backward process, we reconstruct the airfoil by reversing the flow direction. By minimizing the distance between the reconstructed airfoils and actual airfoils, we learn the neural operators that enable us to generate airfoils from functions of Gaussian process. Beyond generation, our method also supports airfoil editing by incorporating the airfoils to be edited as conditions for the generative model.
In summary, our main contributions are threefold: (1) We propose generating airfoil shapes in the functional space to achieve important properties for the aircraft engineering, i.e., arbitrary-resolution control point sampling and maximal design space; (2) We design FuncGenFoil, the first controllable airfoil generative model in the functional space, which effectively incorporates neural operator architectures into the generative model; (3) We further enhance FuncGenFoil with airfoil editing capabilities through minimal adaptations. Experimental results on the AFBench dataset indicate that our proposed method achieves state-of-the-art airfoil generation quality, reducing label error by 74.4% and increasing diversity by 23.2% on the AF-200K dataset, as validated by aerodynamic simulation analysis. In addition, our method is the first to successfully perform airfoil editing by fixing and dragging at any position, achieving nearly zero MSE error (less than $10^{-7}$).
2 Related Works
Generative Models. Generative models based on score matching Ho et al. (2020b); Song et al. (2021) and flow matching Lipman et al. (2023); Tong et al. (2024) have significantly advanced machine learning, achieving state-of-the-art results in areas such as image generation Rombach et al. (2022), text generation Gat et al. (2024), and video generation Ho et al. (2022); Polyak et al. (2024). However, most of these models operate in finite-dimensional spaces and rely on fixed discretizations of data. Such formulations hinder transferability across discretizations and ignore function-level constraints (e.g., continuity, smoothness), motivating the need for function-space generative modeling.
Neural Architecture for Functions space Designing neural architectures to handle function spaces remains a major research challenge. Standard networks typically assume fixed-size inputs, making them unsuitable for arbitrary resolutions. Implicit Neural Representations like SIREN Sitzmann et al. (2020) harness random Fourier features Rahimi and Recht (2007) to represent continuous and differentiable objects through position embeddings. Similarly, NeRF Mildenhall et al. (2021) treats input coordinates as continuous variables, offering flexible resolution for function outputs. Neural operators Anandkumar et al. (2019); Li et al. (2021); Kovachki et al. (2023); Azizzadenesheli et al. (2024) and Galerkin transformers Cao (2021) further generalize neural architectures to process sets of points as functional inputs, enabling function-space learning.
Generative Models in Function Space. In generative modeling, early Neural Processes Garnelo et al. (2018b, a) drew on Gaussian Processes Rasmussen (2003), and later methods such as GASP Dupont et al. (2022b), Functa Dupont et al. (2022a), and GANO Rahman et al. (2022) treat data as function evaluations to enable discretization-independent learning. Energy-based and diffusion models Lim et al. (2023); Kerrigan et al. (2023); Lim et al. (2025); Pidstrigach et al. (2023); Franzese et al. (2024) along with flow-based approaches like FFM Kerrigan et al. (2024) and OpFlow Shi et al. (2024) further extend these ideas. Ultimately, creating comprehensive generative models in function space requires defining suitable stochastic processes, score operators, and consistent neural mappings with specialized training for numerical stability—challenges that remain largely unresolved.
Figure 2: Overview of FuncGenFoil’s neural network. The model is a Fourier Neural Operator, but any other neural operator capable of general-purpose function-space approximation may be used. The model takes as input a function $u_{t}$ (at any resolution $d$), design condition variables $c$, and the generation time $t$. It then processes these as an operator and outputs the current velocity operator $v_{\theta}(u_{t},c,t)$ of the ODE for generation.
Airfoil Design and Optimization. Airfoil design is critical for aerodynamic performance in systems such as aircraft, race cars, and wind turbines. Geometric parameterization enables efficient modeling and optimization of airfoil shapes. Free-Form Deformation (FFD) Mortenson (1999), B-splines Farin (2002), or NURBS Schoenberg (1964)—and is used in CAD/CAE tools like ESP Dannenhoffer (2024) and OpenVSP Hahn (2010)—but its independent control points can lead to local optima, instability, and high dimensionality. Modal parameterization methods, including Proper Orthogonal Decomposition Berkooz et al. (1993), global modal Brüls et al. (2007), and compact modal approaches Li and Zhang (2021), address dimensionality by capturing global features with fewer variables, yet their linear reduction struggles with large deformations and detailed edits. In contrast, the Class-Shape Transformation (CST) Kulfan (2008) offers intuitive physical interpretation and differentiability but is limited in handling significant deformations and is sensitive to parameter selection. Thus, a high-degree-of-freedom geometric representation is needed to support arbitrary shape modifications while accurately capturing highly nonlinear deformations with stability and robustness.
Recently, with the rise of generative models in the AI field, airfoil design has increasingly adopted generative models such as VAE, GAN and diffusion models Chen and Fuge (2021); Li et al. (2022); Xie et al. (2024); Yang et al. (2024); Wei et al. (2024a); Liu et al. (2024), as well as CFD meshes Wei et al. (2023, 2024b), to enhance the representation of airfoil shapes in a generative manner. These models establish an effective transformation from a latent space to the airfoil space. Previous works have primarily focused on modeling the discrete points of airfoils, which limits their flexibility for downstream tasks.
3 FuncGenFoil: Function-space Generative Model for Airfoils
Different from existing airfoil generation methods, FuncGenFoil is constructed as a function-space generative model, capable of producing airfoil geometries as continuous functions instead of discrete points, leveraging the advantages of both parametric-model-based and discrete-point-based methods. We detail the process for airfoil generation and editing tasks, respectively.
3.1 Airfoil Generation
The entire FuncGenFoil is essentially an ordinary differential equation (ODE) generative model that generates airfoil by solving an ODE continuous time $t\in[0,1]$ as follows:
$\mathrm{d}u_{t}=v_{t}\,\mathrm{d}t,$ (1)
where $v_{t}$ is a velocity that gradually transforms an assumed latent coding $u_{0}$ ($u_{t}$ at $t=0$) sampled from stochastic process $\mathcal{P}$ into an airfoil function $u_{1}$ belonging to the target airfoil distribution $\mathcal{Q}$.
Velocity Operator. Obviously, the velocity $v_{t}$ is the key, we achieve it by establishing a parameterized velocity operator $v_{\theta}(u_{t},c,t)$ with model weights $\theta$, i.e., $v_{t}=v_{\theta}(u_{t},c,t)$. To enable the velocity operator $v_{\theta}$ an ability to take the airfoil shape function as input, we build $v_{\theta}$ as a function-space generative model by Operator Flow Matching Shi et al. (2025) as shown in Fig. 2. Specifically, $v_{\theta}$ consists of stacked multiple Fourier Neural Operator layers and it takes noised airfoil $u_{t}$ (at any resolution $d$), condition variables (optional) and a timestamp $t$ as inputs, deriving $v_{t}$.
Figure 3: Airfoil generation by FuncGenFoil. We first sample latent function from a zero-mean Gaussian process with kernel function $K$ at any resolution. Then we solve an ODE flow through $t=0$ to $t=1$ using numerical ODE solver using velocity operator $v_{\theta}$ along this generation process and can generate diverse airfoils. As shown in (c), A single FuncGenFoil model can generate airfoils at arbitrary resolutions, from low to high, while guaranteeing identical outputs when sampling from the same latent function—an inherent property of the model’s architecture.
Airfoil Geometry Parametrization. Since an airfoil has circular topology, we introduce the variable $\alpha\sim[0,2\pi]$ for convenience. We denote $y(\alpha)=u_{1}(\alpha)$ and $x(\alpha)=\frac{\cos(2\pi\alpha)+1}{2}$ [^1] For the remainder of this paper, we omit the explicit dependence on $\alpha$ for clarity., achieving the function of airfoil $u_{0}=(x(\alpha),y(\alpha))$.
Training. We train $v_{\theta}$ under the denoising training process. Given an airfoil geometry $u_{1}$, we compute its corresponding noised sample at time $t$ as follows:
$u_{t}=t\times u_{1}+(1-t)\times u_{0}$ (2)
Then we feed $u_{t}$ into $v_{\theta}$ and train $v_{\theta}$ by matching the velocity operator:
$\mathcal{L}_{\mathrm{FM}}=E_{t\sim[0,1],u_{0},u_{1}}\left[\|v_{t}-v_{\theta}(u_{t},c,t)\|^{2}\right],$ (3)
where $v_{t}$ is the ground-truth velocity:
$v_{t}=u_{1}-u_{0}.$ (4)
Detailed algorithm for training velocity operator is shown in Appendix B Algorithm 2.
Inference. Given a trained velocity operator $v_{\theta}$, the inference process, i.e., airfoil generation process, is equal to deriving airfoil geometry at time $t=1$, denoted as $u_{1}$, based on a latent coding $u_{0}$ sampled from the stochastic process $\mathcal{P}$. $\mathcal{P}$ is assumed as a Gaussian Process $\mathcal{GP}(0,K)$, where $K$ is a covariance kernel function. Therefore, $u_{1}$ could be derived by solving the ODE in Equation 1 numerically (e.g., using the Euler method) as follows:
$u_{1}=u_{0}+\int_{t=0}^{t=1}v_{\theta}(u_{t},c,t)\,\mathrm{d}t.$ (5)
The generation process is shown in Fig. 3 while the detailed algorithm is shown in Appendix B Algorithm 3.
3.2 Airfoil Editing
Figure 4: Airfoil editing by FuncGenFoil. Given a original airfoil $u_{1}$, an editing requirement $\delta$ and target airfoil $u_{1}^{\delta}$. We first infer its latent function $u_{0}$ reversely, and make it learnable as $a_{\theta}$. Then we sample a new airfoil $u_{1}^{{}^{\prime}}$, and conduct a regression in function space via maximum a posteriori estimation $\mathcal{L}_{\mathrm{MAP}}$. After a few iterations of fine-tuning, we can generate edited airfoil $u_{1}^{\delta}$ with high accuracy.
The airfoil editing task enables the user to modify parts of the geometry of a given airfoil, effectively generating a new airfoil geometry while preserving the user-edited sections, denoted as $u_{1}^{\delta}$. We achieve this by an optimization method that maximum a posteriori (MAP) estimation, $\max p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ where $p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ is a probabilistic model that constrains the optimized airfoil $u_{1}^{{}^{\prime}}$ fulfilling the editing requirements and following the generation prior.
To achieve the constraint probability model, we disentangled it with Bayes’ Rule:
$\max_{u_{1}^{{}^{\prime}}}p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})\;=\;\frac{p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})\cdot p(u_{1}^{{}^{\prime}})}{p(u_{1}^{\delta})}$ (6)
$\Rightarrow$ $\max_{u_{1}^{{}^{\prime}}}\log{p(u_{1}^{\delta}|u_{1}^{{}^{\prime}})}+\log{p(u_{1}^{{}^{\prime}})}-\log{p(u_{1}^{\delta})},$
where $p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})$ is a Gaussian measure, so its log term becomes a Mean Square Error (MSE) between $u_{1}^{\delta}$ and $u_{1}^{{}^{\prime}}$, constraining the user edited parts in $u_{1}^{{}^{\prime}}$ keeping consistent with $u_{1}^{\delta}$. $p(u_{1}^{{}^{\prime}})$ is the prior supported by the trained generative model so its log value is derived by Hutchinson trace estimator Hutchinson (1990); Grathwohl et al. (2019) effectively. $p(u_{1}^{\delta})$ is the marginal likelihood, which does not depend on $u_{1}^{{}^{\prime}}$. The final optimization target could be written as:
$\max_{u_{1}^{{}^{\prime}}}\frac{1}{2\sigma^{2}}\sum_{i\in\Delta}(u_{1}^{\delta,i}-u_{1}^{{}^{\prime},i})^{2}+\log{p(u_{1}^{{}^{\prime}})},$ (7)
where $\sigma$ is noise, $\Delta$ denotes point indices in edited part $u_{1}^{\delta}$.
For more realistic generation results, the optimization does not directly adjust $u_{1}^{{}^{\prime}}$; instead, we optimize $u_{1}^{{}^{\prime}}$ indirectly by fine-tuning the entire generative model for a few iterations. The fine-tuning process is illustrated in Fig. 4, while its details are provided in Alg. 1 in the main text. Specifically, we first initialize $u^{\prime}_{1}$ as a resample data of $u_{1}$, by extracting its latent code, denoted as $\alpha_{\theta}$, via the inverse of our generative model and re-generating $u_{1}^{{}^{\prime}}$ from this code. Then we treat Equation 7 as the loss function to train $\theta$ of the velocity operator and $\alpha_{\theta}$ simultaneously. After the model training, the new $u_{1}^{{}^{\prime}}$ is the new edited generation results.
Algorithm 1 Airfoil Editing.
Input: pretrained neural operator $v_{\theta}$, original airfoil function $u_{1}$ (optional) or latent function $u_{0}$ (optional), editing requirement $\delta$, editing resolution $d$, sampling time steps $T$ and steps length $\mathrm{d}t$.
Parameter: gaussian process $\mathcal{GP}(0,K)$ for sampling $u_{0}$, noise level $\sigma$.
Output: new airfoil $\{y_{i}\}$ at resolution $d$.
1:  Inversely sample original airfoil function $\{u_{1,i}\}$ through $v_{\theta}$ and get its latent function $\{u_{0,i}\}$.
2:  Set $\{a_{\theta,i}\}=\{u_{0,i}\}$.
3:  while finetuning… do
4:     Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
5:     Compute $\mathcal{L}_{\mathrm{MAP}}$.
6:     Compute gradient and update $\theta$.
7:  end while
8:  Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
9:  return $\{y_{i}\}=\{u_{1,i}^{{}^{\prime}}\}$
4 Experiments
4.1 Experiment Settings
Tasks. We evaluate two tasks in the airfoil inverse design problem: conditional airfoil generation and freestyle airfoil editing. In the conditional generation task, the model is given a set of 11 geometric parameters that describe the airfoil geometry. Detailed parameter definitions are provided in the Supplementary Materials. The model must generate airfoils that satisfy these geometric constraints. In the freestyle editing task, the model takes an original airfoil and a target modification, such as adjusting the position of a specific point on the airfoil. The selected point can be any location on the airfoil surface. The model must generate an airfoil that reflects the specified modification.
Metrics. We adopt the metrics introduced in AFBench Liu et al. (2024) to evaluate the generated airfoils:
Label Error measures the difference between the physical parameters of the generated or edited airfoil and the intended target parameters, calculated as $\sigma_{i}=\left|\hat{p}_{i}-p_{i}\right|$, where $\sigma_{i}$ is the label error for the $i$-th parameter, $\hat{p}_{i}$ is the $i$-th geometric parameter from the generated airfoil, and $p_{i}$ is the corresponding target parameter. Smaller values indicate better alignment with the target parameters.
Diversity quantifies the variety of generated airfoils, calculated as $D=\frac{1}{n}\sum_{i=1}^{n}\log\det(L_{S_{i}})$, where $n$ is the number of samples, and $L_{S_{i}}$ is the similarity matrix of the $i$-th subset, computed based on Euclidean distances between airfoils in the subset. Higher values indicate greater diversity among generated airfoils.
Smoothness measures the geometric smoothness of the generated airfoils, calculated as:
$\small M=\sum_{i=1}^{N}\text{Distance}(P_{n},P_{n-1}P_{n+1}),$ (8)
where $P_{n}$ is the $n$-th point, and $P_{n-1}P_{n+1}$ is the line segment between adjacent points. The function $\text{Distance}(P_{n},P_{n-1}P_{n+1})$ computes the perpendicular distance from $P_{n}$ to this line segment. Smaller values indicate better geometric quality.
Datasets. To benchmark our method, we conduct experiments on three datasets: UIUC Selig (1996), Supercritical Airfoil (Super), and AF-200K. UIUC contains 1,600 designed airfoil geometries. Super focuses on supercritical airfoils and includes approximately 20,000 airfoil samples. AF-200K consists of 200,000 highly diversified airfoil samples.
Baselines. We include the baseline models proposed in AFBench Liu et al. (2024), specifically the conditional VAE (CVAE), conditional GAN (CGAN), the modified VAE with PARSEC parameters and control keypoints (PK-VAE), along with PK-GAN, PKVAE-GAN, the U-Net-based PK-DIFF, and the transformer-based PK-DIT.
Table 1: Quantitative comparison between FuncGenFoil and baseline methods across different datasets for conditional generation task. Label error, diversity, and smoothness of the generated airfoils are reported.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
CVAE AF-200K 72.9 52.5 35.2 15900 99 95 29000 19.1 15.3 46 104 4131 -155.4 7.09
CGAN AF-200K 107 85.0 54.4 23200 143 137 59600 25.3 22.3 53 129 7596 -120.5 7.31
PK-VAE AF-200K 63.0 47.9 31.3 8620 66 64 17100 13.5 9.3 33 78 2375 -150.1 5.93
PK-GAN AF-200K 81.8 63.0 47.0 21030 120 117 32470 22.5 19.6 50 122 4923 -112.3 3.98
PKVAE-GAN AF-200K 56.8 31.7 31.0 5650 46 43 12000 9.1 5.1 28 63 1633 -129.6 2.89
PK-DIT AF-200K 11.2 32.3 15.4 1050 13 11.5 9790 0.5 0.5 23 24 997 -93.2 1.04
FuncGenFoil AF-200K 1.84 17.4 0.56 721 47.7 0.98 1676 0.45 0.65 160 174 255 -71.6 1.41
PK-DIT UIUC 63.8 51.4 33.6 11830 87 84.9 25700 16.9 11.9 36 98 3456 -141.7 6.03
FuncGenFoil UIUC 14.2 12.7 0.86 434.9 90.7 3.7 3559 0.55 0.69 129.9 135.2 398.4 -66.3 1.17
PK-DIT Super 52.0 35.0 24.0 3010 29 33.2 10500 8.3 2.6 27 33 1250 -123.4 1.97
FuncGenFoil Super 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
Table 2: Quantitative evaluation of the airfoil editing task across different edit scales. The mean squared error (MSE) between the generated and target airfoils as well as the smoothness of the generated airfoils are reported.
[…]
Super 0.0001 2.41 1.16
0.0002 2.45 1.15
0.0004 2.75 1.15
Table 3: Quantitative evaluation of our model (training resolution=257) across different sampling resolutions for conditional generation task.
[…]
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
Super 257 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
513 0.74 2.59 0.13 55.1 1.54 0.09 57.7 0.07 0.08 41.1 40.0 18.1 -100.5 0.49
1025 0.75 3.18 0.17 56.5 2.03 0.12 59.2 0.07 0.09 60.8 55.0 21.6 -92.9 0.31
Implementation Details. On the AF-200K dataset, we trained for 2 million iterations with a batch size of 2,048 using 8 NVIDIA 4090 GPUs. On the Supercritical Airfoil dataset, we trained for 1 million iterations with a batch size of 256 on a single NVIDIA 4090 GPU. We use the Matérn kernel function as the kernel for the Gaussian process at time $t=0$. Other training hyperparameters are detailed in the Supplementary Materials.
4.2 Main Results
Conditional Airfoil Generation.
As shown in Table 1, FuncGenFoil significantly outperforms the strongest baseline method (PK-DIT) across all metrics. For label error, FuncGenFoil achieves a reduction in average error by 88.5% on UIUC, 99% on Super, and 74.4% on AF-200K. This substantial decrease in label error underscores the effectiveness of our approach in generating airfoils that more precisely adhere to target geometric parameters. In terms of diversity, FuncGenFoil demonstrates notable improvements, surpassing the best baseline methods by 75.4, 14.6, and 21.6 on UIUC, Super, and AF-200K, respectively. This highlights the model’s superior capability in capturing and generating a broader spectrum of valid airfoil designs. Additionally, our generated airfoils exhibit enhanced surface smoothness, as evidenced by reductions of 0.99 and 4.86 in smoothness values ($10^{-2}$) on the UIUC and Super datasets, respectively. This improvement is particularly crucial for aerodynamic performance, as smoother airfoil surfaces contribute to reduced drag and improved flow characteristics.
Freestyle Airfoil Editing. Table 2 shows the average performance across 300 editing cases using FuncGenFoil. In each case, the model adjusts 2 to 4 randomly selected positions on the wing surface to target locations over 10 fine-tuning steps, with edit scales ranging from $1\times 10^{-4}$ to $4\times 10^{-4}$. The results demonstrate that FuncGenFoil can perform accurate airfoil editing with minimal errors (less than $2.75\times 10^{-7}$ MSE) and high surface smoothness (less than 1.16 ×$\times 10^{-2}$ smoothness value) with just a few fine-tuning steps. Fig. 5 illustrates the editing requirements and generated airfoils during the fine-tuning stage. It shows that our model faithfully completes the freestyle editing task by generating reasonable airfoils according to user-specified editing requirements.
Figure 5: Examples of two FuncGenFoil instances performing airfoil editing over 100 training iterations. The red section represents the editing requirement, while the black airfoil shows the generated samples at each training stage. The results demonstrate that the generated airfoil quickly adapts to the editing requirements within a few iterations, achieving a natural and smooth function regression.
Any-Resolution Airfoil Generation. One of the key advantages of FuncGenFoil is its ability to generate airfoils at any resolution while maintaining high generation quality. This is achieved by representing airfoils as functions and learning resolution-independent function transformations, enabling flexible and consistent airfoil generation across different scales. To evaluate high-resolution generation, we use the model trained on the Supercritical Airfoil dataset at a resolution of 257 and sample new airfoils at resolutions of 513 and 1025. We then assess these higher-resolution airfoils using the same metrics as in the conditional generation task. The results are presented in Table 3. We observe consistent generation quality at $\times$ and even $\times$ the training resolution, with a maximum increase of 9.1×$\times 10^{-3}$ in average label error, a decrease of 16.1 in diversity, and an increase of 0.67×$\times 10^{-2}$ in smoothness value.
Aerodynamic Simulation. To further assess the physical properties of the generated airfoils and validate the effectiveness of our method, we analyze their aerodynamic performance using the NASA Common Research Model (CRM)  [^2] https://commonresearchmodel.larc.nasa.gov/ and perform Reynolds-Averaged Navier-Stokes (RANS) computational fluid dynamics (CFD) simulations on the generated samples. The CRM dataset contains 135,000 CRM wing geometries along with their corresponding aerodynamic performance, computed using the RANS CFD solver ADflow Mader et al. (2020). We pretrain the FuncGenFoil model on all 135,000 CRM wing geometries and generate 500 new CRM wing geometries for RANS CFD evaluation. We analyze the lift-to-drag ratio ($L/D$) of these newly generated geometries and compare them with the original dataset, as shown in Fig. 6. Our results indicate that the $L/D$ distribution of the generated samples closely aligns with that of the original dataset, with the highest density occurring around $L/D=19$, demonstrating the model’s ability to learn and generate physically plausible wing geometries. Additionally, we visualize the coefficient of pressure contours for selected CFD cases in Fig. 7. These visualizations confirm that FuncGenFoil can generate airfoils with diverse aerodynamic performance characteristics.
Figure 6: Comparison of the lift-to-drag ratio histograms for CRM wings in the original dataset and the samples generated by our FuncGenFoil model.
Figure 7: Visualization of the pressure coefficient for the generated CRM wings, obtained through aerodynamic simulation.
4.3 Ablation Study
We first ablate the use of the Matérn kernel $K$ with different parameters as the kernel function for the sampled Gaussian process. As shown in Table 4, we compare the cases of Matérn kernel $\nu=1.5$, $l=0.01$, and Matérn kernel $\nu=2.5$, $l=0.03$, where the latter is smoother due to the combined effect of both parameters. Notably, the limits of $l=0$ and $\nu\to\infty$ correspond to pure white noise and the RBF kernel, respectively. To further highlight the importance of kernel function selection, we also test two models of different sizes (by setting different FNO modes). From Table 4, we observe that a smoother kernel improves both the smoothness and accuracy of the generated airfoils. Additionally, the comparison highlights that, under our setup, selecting a smoother kernel leads to a more substantial performance improvement than increasing the model size. This underscores the benefit of modeling the function object in the airfoil generation task rather than modeling discrete points.
In Table 5, we compare the MSE of constraint condition calculations in the airfoil editing task using two initialization schemes: (1) using the latent variables obtained through ODE inversion of the original airfoil as the prior, and (2) using a zero prior $u_{0}=0$ without ODE inversion. Both schemes are fine-tuned with the same number of iterations. As shown, the ODE inversion-based prior significantly reduces the MSE compared to the zero prior, with a maximum reduction of 84.85. This underscores the importance of incorporating original airfoil information during the initialization process.
Table 4: Ablation study on the impact of different kernels and model sizes for the conditional airfoil generation task.
Matérn $\nu$ Matérn $l$ FNO modes $\bar{\sigma}\downarrow$ $\mathcal{M}\downarrow$
1.5 0.01 32 0.108 0.0119
2.5 0.03 32 0.013 0.0098
1.5 0.01 48 0.039 0.0101
2.5 0.03 48 0.013 0.0098
Table 5: Ablation study on the effect of different latent variable initialization methods for the airfoil editing task.
Edit scale MSE$\mathcal{\downarrow}(10^{-7})$
w/ ODE inversion w/o ODE inversion
0.0001 2.42 ($\downarrow$83.88) 86.3
0.0002 2.46 ($\downarrow$84.14) 86.6
0.0004 2.75 ($\downarrow$84.85) 87.6
5 Conclusion
In this work, we tackle the critical challenge of generating high-fidelity airfoil geometries that effectively balance expressiveness, smoothness, and resolution flexibility. We introduce FuncGenFoil, a function-space generative model leveraging neural operators and flow matching, which represents airfoils as continuous, smooth geometries without resolution constraints while preserving the expressiveness of data-driven methods. Comprehensive experimental results show that FuncGenFoil outperforms state-of-the-art techniques in terms of label error, diversity, and smoothness, highlighting its potential for high-fidelity airfoil design. Furthermore, the generated wing geometries have been validated through aerodynamic simulations. This work paves the way for more efficient, scalable, and versatile airfoil generation, with significant applications in aerodynamic shape optimization for aircraft manufacturing.
Appendix A Airfoil Geometry Parameters
Table 6: Geometric parameters of the airfoils used in conditional generation tasks
Index Symbol Meaning
1 $R_{\mathrm{le}}$ the leading edge radius
2 $X_{\mathrm{up}}$ upper crest position x
3 $Y_{\mathrm{up}}$ upper crest position y
4 $Z_{\mathrm{xxup}}$ upper crest curvature
5 $X_{\mathrm{lo}}$ lower crest position x
6 $Y_{\mathrm{lo}}$ lower crest position y
7 $Z_{\mathrm{xxlo}}$ lower crest curvature
8 $Y_{\mathrm{te}}$ trailing edge position
9 $\Delta Y_{\mathrm{te}}$ trailing thickness
10 $\alpha_{\mathrm{te}}$ trailing edge angle up
11 $\beta_{\mathrm{te}}$ trailing edge angle down
Algorithm 2 Generative Model Training.
Input: data resolution $d$, data $u_{1}$, design condition variables $c$ (optional).
Parameter: gaussian process $\mathcal{GP}(0,K)$ for sampling $u_{0}$.
Output: velocity operator $v_{\theta}$.
1:  while training… do
2:     sample $t\sim[0,1]$, $u_{0}\sim\mathcal{GP}(0,K)$ at resolution $d$ and get $\{u_{0,i}\}$.
3:     Compute $v_{t}$ at resolution $d$ and get $\{v_{t,i}\}$.
4:     Compute $u_{t}$ at resolution $d$ and get $\{u_{t,i}\}$.
5:     Compute $v_{\theta}(u_{t},c,t)$ at resolution $d$ and get $\{v_{\theta,i}\}$.
6:     Minimize $\|\{v_{\theta,i}\}-\{v_{t,i}\}\|^{2}$.
7:     Compute gradient and update $\theta$.
8:  end while
9:  return $v_{\theta}$.
Algorithm 3 Inference Process (Airfoil Generation)
Input: sampling resolution $d$, sampling time steps $T$ and steps length $\mathrm{d}t$, latent function $u_{0}$ (optional), design condition variables $c$ (optional).
Parameter: gaussian process $\mathcal{GP}(0,K)$ for sampling $u_{0}$.
Output: airfoil $\{y_{i}\}$ at resolution $d$.
1:  Let $t=0$, $u_{0}\sim\mathcal{GP}(0,K)$ at resolution $d$ and get $\{u_{0,i}\}$.
2:  while $t<=1$ do
3:     Compute $v_{\theta}(u_{t},c,t)$ at resolution $d$ and get $\{v_{\theta,i}\}$.
4:     Compute $\{u_{t+\mathrm{d}t,i}\}=\{u_{t,i}\}+\{v_{\theta,i}\mathrm{d}t\}$.
5:     $t=t+\mathrm{d}t$.
6:  end while
7:  return $\{y_{i}\}=\{u_{1,i}\}$
Table 7: Training hyperparameters and model parameters in conditional airfoils generation tasks
Training hyperparameters Value
max learning rate 5e-6
batch size 512
optimizer used in pre-training and finetuning Adam
train max iterations 1,000,000
ODE Solver time steps 100
Model parameters Value
Fourier neural operator layers 8
Fourier neural operator modes 48
Fourier neural operator hidden channels 256
Matérn kernel $\nu$ 2.5
Matérn kernel $l$ 0.03
We adopted the same conditions and symbols as those in AFBench. The parameters below, including the leading-edge radius, the curvature at the maximum thickness of the upper and lower surfaces, and the trailing-edge angle of the airfoil, were obtained through fitting. Note that our implementation of the angle calculation differs slightly from that of AFBench, as we found that the original fitting method was not invariant when the sampling precision increased.
Appendix B Algorithm Details
The algorithm for training airfoil generative models is shown in Algorithm 2, and inference using a trained model is shown in Algorithm 3. In the actual implementation, we utilized Optimal Transport techniques and employed the Matérn kernel as the kernel function for the Gaussian process in the latent space of the generative model.
Appendix C Training Details
In most of the experiments, we followed the hyperparameters listed in the table below to train the models, including both the pre-training and fine-tuning stages. On the AF-200K dataset, we increased the maximum number of iterations to 2,000,000 and the batch size to 2048. For the number of time steps used to solve the ODE, we tested reconstruction metrics on the super dataset with 10, 100, and 1000 steps, and ultimately chose 100 steps as the default setting for all other experiments, except for the fine-tuning stage, where we used 10 steps for training efficiency.

### v3

@@
 Abstract
-Aircraft manufacturing is the jewel in the crown of industry, among which generating high-fidelity airfoil geometries with controllable and editable representations remains a fundamental challenge.
-While existing deep-learning-based methods rely on predefined parametric function families, e.g., Bézier curves and discrete point-based representations, they suffer from inherent trade-offs between expressiveness and resolution flexibility.
-To tackle this challenge, we introduce FuncGenFoil, a novel function-space generative model that directly learns functional airfoil geometries.
-Our method inherits both the advantages of arbitrary resolution sampling and the smoothness of parametric functions, as well as the strong expressiveness of discrete point-based functions.
-Empirical evaluations on the AFBench dataset demonstrate that FuncGenFoil improves upon state-of-the-art methods in airfoil generation by achieving a relative -74.4% label error reduction and +23.2% diversity increase on the AF-200K dataset.
+Aircraft manufacturing is the jewel in the crown of industry, in which generating high-fidelity airfoil geometries with controllable and editable representations remains a fundamental challenge.
+Existing deep learning methods, which typically rely on predefined parametric representations (e.g., Bézier curves) or discrete point sets, face an inherent trade-off between expressive power and resolution adaptability.
+To tackle this challenge, we introduce FuncGenFoil, a novel function-space generative model that directly reconstructs airfoil geometries as function curves.
+Our method inherits the advantages of arbitrary-resolution sampling and smoothness from parametric functions, as well as the strong expressiveness of discrete point-based representations.
+Empirical evaluations demonstrate that FuncGenFoil improves upon state-of-the-art methods in airfoil generation, achieving a relative 74.4% reduction in label error and a 23.2% increase in diversity on the AF-200K dataset.
 Our results highlight the advantages of function-space modeling for aerodynamic shape optimization, offering a powerful and flexible framework for high-fidelity airfoil design.
-Our code will be released.
 1 Introduction
+The airfoil inverse design problem serves as a central aspect of aircraft manufacturing.
+Traditionally, given geometric requirements, engineers first select the most similar airfoils from well-known airfoil datasets (e.g., NACA [8]) and rely on a trial-and-error strategy [51].
+Considering the mission of the aircraft, an initial airfoil design that meets these conditions is preliminarily created.
+Then, through iterative rounds of physical analyses, such as aerodynamic and mechanical evaluations, the airfoil is optimized to achieve improved performance until the specified requirements are met.
+In practice, such direct design procedures are highly inefficient and time-consuming, often taking months.
+To minimize development and design time, as well as associated costs, automated design methods have been introduced as efficient alternatives in aircraft manufacturing engineering.
+In particular, machine learning-based design and optimization techniques have gained significant attention.
+However, before applying these algorithms to airfoil design, it is crucial to determine appropriate methods for representing airfoils within these algorithms.
 Figure 1: The conceptual difference between FuncGenFoil and previous airfoil representation methods.
@@
 In contrast, (c) illustrates FuncGenFoil’s approach, where an airfoil is treated as a continuous function mapped from a latent function, enabling a generative model in function space.
-The airfoil inverse design problem serves as a central aspect of aircraft manufacturing.
-Traditionally, given geometric requirements, engineers first select the most similar airfoils from well-known airfoil datasets (e.g., NACA Cummings et al.
-(2015)) and leverage the trial-and-error strategy Sharma et al.
-(2021).
-Considering the mission of the aircraft, an initial airfoil design that meets the design conditions is preliminarily created.
-Then, through rounds of physical analysis, such as aerodynamics and mechanics, the airfoil is iteratively optimized to achieve better performance until the requirements are met.
-In practice, such direct design procedures are highly inefficient and time-consuming, often taking months.
-To minimize development and design time, as well as associated costs, automatic design methods have been introduced as efficient alternatives in aircraft manufacturing engineering.
-In particular, machine learning-based design and optimization techniques have gained significant attention.
-However, before applying algorithms to airfoil design, it is crucial to determine appropriate methods for representing airfoils within these algorithms.
-Existing methods for airfoil representation can generally be divided into two categories: parametric-model-based approaches Xie et al.
-(2024) and discrete-point-based methods Liu et al.
-(2024); Sekar et al.
-(2019).
-First, parametric-model-based methods predefine function families, e.g., Bézier curves Chen and Fuge (2021), Hicks-Henne curves, and NURBS, and leverage mathematical optimization techniques or generative models to determine the parameters of these functions for a new airfoil.
-These methods rigorously preserve key geometric properties of the defined functional families, e.g., high-order smoothness.
-Furthermore, such functional representations of airfoils allow for arbitrary sampling of control points in real manufacturing, given the constraints of engineering precision.
-Despite these benefits, parametric-model-based methods suffer from a significantly reduced design space, i.e., selecting a specific function family excludes the possibility of other shapes, which limits the upper bound of airfoil design algorithms.
+Existing methods for airfoil representation can generally be divided into two categories: parametric-model-based approaches [61] and discrete-point-based methods [37, 49].
+First, parametric-model-based methods predefine function families, e.g., Bézier curves [7], Hicks-Henne curves, and NURBS, and leverage mathematical optimization techniques or generative models to determine the parameters of these functions for generating a new airfoil.
+These methods rigorously preserve key geometric properties of the defined function families, e.g., higher-order smoothness.
+Furthermore, such functional representations of airfoils allow arbitrary sampling of control points in real manufacturing, given engineering precision constraints.
+Despite these benefits, parametric-model-based methods suffer from a significantly reduced design space; i.e., selecting a specific function family excludes other possible shapes, thus limiting the upper bound of airfoil design algorithms.
 Second, discrete-point-based methods directly generate multiple points to represent airfoil shapes.
-These methods maximize the design space of airfoils but cannot maintain some important mathematical properties, e.g., continuity.
-Furthermore, they cannot directly generate control points at arbitrary resolutions because the number of generated points is typically fixed for each model after training.
+These methods maximize the airfoil design space but cannot inherently maintain certain important mathematical properties, e.g., continuity.
+Furthermore, they cannot directly generate control points at arbitrary resolutions, as the number of generated points is typically fixed for each model after training.
 To address the trade-offs between these two mainstream approaches, we ask: Can we design an algorithm that leverages the advantages of both?
-In this paper, we answer this question by proposing FuncGenFoil, a novel function-space generative model for airfoil representation (see Fig.
-1).
-Different from previous data-driven and generative methods, e.g., cVAE Kingma (2013) and cGAN Mirza (2014), which directly generate discrete points, our approach models airfoil geometry using a general continuous function approximator Anandkumar et al.
-(2019); Li et al.
-(2021); Kovachki et al.
-(2023); Azizzadenesheli et al.
-(2024), i.e., neural operator architectures.
-Simultaneously, it leverages recent advances in generative methods, e.g., diffusion models Ho et al.
-(2020a) and flow matching models Lipman et al.
-(2022), to generate diverse airfoils beyond the design space of any predefined geometry function.
-Our method has the advantages of both parametric-model-based approaches and discrete-point-based methods.
-Due to its functional representation, the generated airfoil is continuous and can be sampled at arbitrary resolutions, making it easier to manufacture.
-Moreover, thanks to the general functional approximator nature of neural operators, our method can explore a broader design space beyond predefined functional families.
-Specifically, we leverage the flow matching framework Lipman et al.
-(2022), an improved alternative to diffusion models, and FNO Li et al.
-(2021), a resolution-free neural operator, as the backbone of our generative model to design FuncGenFoil.
-In the forward process, we perturb the airfoil into a noise distribution through straight flows.
-In the backward process, we reconstruct the airfoil by reversing the flow direction.
-By minimizing the distance between the reconstructed airfoils and actual airfoils, we learn the neural operators that enable us to generate airfoils from functions of Gaussian process.
-Beyond generation, our method also supports airfoil editing by incorporating the airfoils to be edited as conditions for the generative model.
-In summary, our main contributions are threefold: (1) We propose generating airfoil shapes in the functional space to achieve important properties for the aircraft engineering, i.e., arbitrary-resolution control point sampling and maximal design space; (2) We design FuncGenFoil, the first controllable airfoil generative model in the functional space, which effectively incorporates neural operator architectures into the generative model; (3) We further enhance FuncGenFoil with airfoil editing capabilities through minimal adaptations.
-Experimental results on the AFBench dataset indicate that our proposed method achieves state-of-the-art airfoil generation quality, reducing label error by 74.4% and increasing diversity by 23.2% on the AF-200K dataset, as validated by aerodynamic simulation analysis.
-In addition, our method is the first to successfully perform airfoil editing by fixing and dragging at any position, achieving nearly zero MSE error (less than $10^{-7}$).
+In this paper, we address this question by proposing FuncGenFoil, a novel function-space generative model for airfoil representation (see Figure 1).
+Unlike previous data-driven generative methods in finite-dimensional spaces, e.g., cVAE [25], cGAN [41], diffusion models [19], or flow matching models [35], which directly generate discrete point outputs, we leverage recent advances in diffusion and flow models defined in function space [33, 23, 32, 42, 13, 52] to produce infinite-dimensional outputs as general continuous functions.
+Simultaneously, our approach models airfoil geometry using Neural Operator architectures [1, 31, 27, 2], enabling the generation of diverse airfoils beyond the design spaces of predefined geometric function families, thanks to its general operator approximation capability.
+Our method combines the advantages of both parametric-model-based and discrete-point-based approaches.
+Due to its intrinsic representation in function space, the generated airfoils are continuous, and both training and sampling can be performed at arbitrary resolutions, facilitating downstream optimization and manufacturing processes.
+Specifically, we use the flow matching framework [35], an improved alternative to diffusion models, and FNO [31], a resolution-free neural operator, as the backbone of our generative model to design FuncGenFoil.
+During training, we perturb the airfoil into a noise function through straight flows and let neural operators learn the airfoil’s deconstructed velocity via flow matching.
+During inference, we reconstruct the airfoil by reversing the flow direction starting from a Gaussian process.
+Beyond generation, our method also supports airfoil editing, generating new airfoils from an original airfoil by conforming to certain geometric constraints.
+This can be accomplished by a few steps of fine-tuning with no additional data other than the original airfoil.
+Constraints can be provided in various forms, such as specifying contour points that the airfoil must pass through or setting geometric constraint parameters, such as thickness.
+In summary, our main contributions are threefold: (1) We propose generating airfoil shapes in function space to achieve important properties for aircraft engineering, i.e., arbitrary-resolution control point sampling and maximal design space; (2) We design FuncGenFoil, the first controllable airfoil generative model in function space, which effectively incorporates neural operator architectures into the generative modeling framework; and (3) We further enhance FuncGenFoil with airfoil editing capabilities through minimal adaptations.
+Experimental results indicate that our proposed method achieves state-of-the-art airfoil generation quality, reducing label error by 74.4% and increasing diversity by 23.2% on the AF-200K dataset, as validated through aerodynamic simulation analysis.
+In addition, our method is the first to successfully perform airfoil editing by fixing or dragging points at arbitrary positions, achieving nearly zero MSE (less than $10^{-7}$).
 2 Related Works
 Generative Models.
-Generative models based on score matching Ho et al.
-(2020b); Song et al.
-(2021) and flow matching Lipman et al.
-(2023); Tong et al.
-(2024) have significantly advanced machine learning, achieving state-of-the-art results in areas such as image generation Rombach et al.
-(2022), text generation Gat et al.
-(2024), and video generation Ho et al.
-(2022); Polyak et al.
-(2024).
-However, most of these models operate in finite-dimensional spaces and rely on fixed discretizations of data.
-Such formulations hinder transferability across discretizations and ignore function-level constraints (e.g., continuity, smoothness), motivating the need for function-space generative modeling.
-Neural Architecture for Functions space Designing neural architectures to handle function spaces remains a major research challenge.
+Generative models based on score matching [20, 55] and flow matching [34, 57, 36] have significantly advanced machine learning, achieving state-of-the-art results in areas such as image generation [47], text generation [16], and video generation [21, 43].
+However, most of these models operate in finite-dimensional spaces and rely on fixed discretizations of the data.
+Such formulations hinder transferability across discretizations and neglect function-level constraints (e.g., continuity, smoothness), motivating the need for generative modeling in function space.
+Neural Architectures for Function Modeling.
+Designing neural architectures to handle function spaces remains a major research challenge.
 Standard networks typically assume fixed-size inputs, making them unsuitable for arbitrary resolutions.
-Implicit Neural Representations like SIREN Sitzmann et al.
-(2020) harness random Fourier features Rahimi and Recht (2007) to represent continuous and differentiable objects through position embeddings.
-Similarly, NeRF Mildenhall et al.
-(2021) treats input coordinates as continuous variables, offering flexible resolution for function outputs.
-Neural operators Anandkumar et al.
-(2019); Li et al.
-(2021); Kovachki et al.
-(2023); Azizzadenesheli et al.
-(2024) and Galerkin transformers Cao (2021) further generalize neural architectures to process sets of points as functional inputs, enabling function-space learning.
+Implicit neural representations, such as SIREN [54], harness random Fourier features [44] to represent continuous and differentiable objects through position embeddings.
+Similarly, NeRF [40] treats input coordinates as continuous variables, offering flexible resolution for function outputs.
+Neural operators [1, 31, 27, 2] and Galerkin transformers [5] further generalize neural architectures to process sets of points as functional inputs, enabling function-space learning.
+Recently, a neural operator with localized integral and differential kernels has been introduced to improve its capability in capturing local features [38].
+To handle complex and variable geometries, the point cloud neural operator has been proposed, which adaptively processes point cloud datasets [63].
 Generative Models in Function Space.
-In generative modeling, early Neural Processes Garnelo et al.
-(2018b, a) drew on Gaussian Processes Rasmussen (2003), and later methods such as GASP Dupont et al.
-(2022b), Functa Dupont et al.
-(2022a), and GANO Rahman et al.
-(2022) treat data as function evaluations to enable discretization-independent learning.
-Energy-based and diffusion models Lim et al.
-(2023); Kerrigan et al.
-(2023); Lim et al.
-(2025); Pidstrigach et al.
-(2023); Franzese et al.
-(2024) along with flow-based approaches like FFM Kerrigan et al.
-(2024) and OpFlow Shi et al.
-(2024) further extend these ideas.
-Ultimately, creating comprehensive generative models in function space requires defining suitable stochastic processes, score operators, and consistent neural mappings with specialized training for numerical stability—challenges that remain largely unresolved.
-Figure 2: Overview of FuncGenFoil’s neural network.
-The model is a Fourier Neural Operator, but any other neural operator capable of general-purpose function-space approximation may be used.
-The model takes as input a function $u_{t}$ (at any resolution $d$), design condition variables $c$, and the generation time $t$.
-It then processes these as an operator and outputs the current velocity operator $v_{\theta}(u_{t},c,t)$ of the ODE for generation.
+Early Neural Processes [15, 14] drew upon Gaussian processes [46], and later methods such as GASP [11], Functa [10], and GANO [45] treat data as function evaluations to enable discretization-independent learning.
+Energy-based and diffusion models [33, 23, 32, 42, 13], along with flow-based approaches like FFM [24] and OpFlow [52], further extend these ideas.
+Ultimately, developing comprehensive generative models in function space requires defining suitable stochastic processes, score operators, and consistent neural mappings, along with specialized training methods for numerical stability—challenges that remain largely unresolved.
 Airfoil Design and Optimization.
-Airfoil design is critical for aerodynamic performance in systems such as aircraft, race cars, and wind turbines.
-Geometric parameterization enables efficient modeling and optimization of airfoil shapes.
-Free-Form Deformation (FFD) Mortenson (1999), B-splines Farin (2002), or NURBS Schoenberg (1964)—and is used in CAD/CAE tools like ESP Dannenhoffer (2024) and OpenVSP Hahn (2010)—but its independent control points can lead to local optima, instability, and high dimensionality.
-Modal parameterization methods, including Proper Orthogonal Decomposition Berkooz et al.
-(1993), global modal Brüls et al.
-(2007), and compact modal approaches Li and Zhang (2021), address dimensionality by capturing global features with fewer variables, yet their linear reduction struggles with large deformations and detailed edits.
-In contrast, the Class-Shape Transformation (CST) Kulfan (2008) offers intuitive physical interpretation and differentiability but is limited in handling significant deformations and is sensitive to parameter selection.
-Thus, a high-degree-of-freedom geometric representation is needed to support arbitrary shape modifications while accurately capturing highly nonlinear deformations with stability and robustness.
-Recently, with the rise of generative models in the AI field, airfoil design has increasingly adopted generative models such as VAE, GAN and diffusion models Chen and Fuge (2021); Li et al.
-(2022); Xie et al.
-(2024); Yang et al.
-(2024); Wei et al.
-(2024a); Liu et al.
-(2024), as well as CFD meshes Wei et al.
-(2023, 2024b), to enhance the representation of airfoil shapes in a generative manner.
-These models establish an effective transformation from a latent space to the airfoil space.
-Previous works have primarily focused on modeling the discrete points of airfoils, which limits their flexibility for downstream tasks.
+Airfoil design is essential for aircraft and wind turbines.
+Geometric parameterization supports efficient shape modeling and optimization.
+Techniques such as Free-Form Deformation (FFD) [12] and NURBS [48] are widely used in CAD/CAE tools [9, 18], but their independent control points can cause instability, high dimensionality, and suboptimal solutions.
+Modal methods, including Proper Orthogonal Decomposition [3], global modal [4], and compact modal [30] approaches, reduce dimensionality by encoding global features, though they struggle with large or detailed shape changes.
+Class-Shape Transformation (CST) [28] offers interpretable and differentiable shapes but lacks flexibility for large deformations and is sensitive to parameter choices.
+Therefore, a high-degree-of-freedom representation is needed for robust, nonlinear shape modeling.
+Unlike general computer vision tasks [56, 26], engineering AI models are domain-specific and often limited by sparse data.
+Generative models are advantageous here, as they can exploit unlabeled or untested samples.
+In airfoil design, VAEs, GANs, and diffusion models [7, 29, 61, 62, 58, 37], along with CFD-based mesh representations [59, 60], have been used to map latent spaces to airfoil shapes.
+However, most existing methods model discrete airfoil points, limiting flexibility for downstream applications.
 3 FuncGenFoil: Function-Space Generative Model for Airfoils
-Different from existing airfoil generation methods, FuncGenFoil is constructed as a function-space generative model, capable of producing airfoil geometries as continuous functions instead of discrete points, leveraging the advantages of both parametric-model-based and discrete-point-based methods.
-We detail the process for airfoil generation and editing tasks, respectively.
+In contrast to existing airfoil generation methods, FuncGenFoil is constructed as a function-space generative model capable of producing airfoil geometries as continuous functions rather than discrete points, thereby leveraging the advantages of both parametric-model-based and discrete-point-based methods.
+In this section, we detail the processes for airfoil generation and editing tasks, respectively.
 3.1 Airfoil Generation
-The entire FuncGenFoil is essentially an ordinary differential equation (ODE) generative model that generates airfoil by solving an ODE continuous time $t\in[0,1]$ as follows:
-$\mathrm{d}u_{t}=v_{t}\,\mathrm{d}t,$ (1)
-where $v_{t}$ is a velocity that gradually transforms an assumed latent coding $u_{0}$ ($u_{t}$ at $t=0$) sampled from stochastic process $\mathcal{P}$ into an airfoil function $u_{1}$ belonging to the target airfoil distribution $\mathcal{Q}$.
+Airfoil Parametrization.
+Since an airfoil curve $(x,y)$ has circular topology, we introduce the variable $\alpha\in[0,1]$ as the domain of the function, denoting $y(\alpha)=f(x(\alpha))$ and $x(\alpha)=\frac{\cos(2\pi\alpha)+1}{2}$, [^1] For the remainder of this paper, we omit the explicit dependence on $\alpha$ for clarity and convenience.
+as shown in Figure 7.
+The entire FuncGenFoil framework is essentially an ordinary differential equation (ODE)-based generative model that produces airfoil curve functions by solving an ODE along continuous time $t\in[0,1]$ as $u_{1}=u_{0}+\int_{t=0}^{t=1}v_{t}\,\mathrm{d}t$, where $u_{1}=y$, starting from some latent function $u_{0}$.
 Velocity Operator.
-Obviously, the velocity $v_{t}$ is the key, we achieve it by establishing a parameterized velocity operator $v_{\theta}(u_{t},c,t)$ with model weights $\theta$, i.e., $v_{t}=v_{\theta}(u_{t},c,t)$.
-To enable the velocity operator $v_{\theta}$ an ability to take the airfoil shape function as input, we build $v_{\theta}$ as a function-space generative model by Operator Flow Matching Shi et al.
-(2025) as shown in Fig.
-2.
-Specifically, $v_{\theta}$ consists of stacked multiple Fourier Neural Operator layers and it takes noised airfoil $u_{t}$ (at any resolution $d$), condition variables (optional) and a timestamp $t$ as inputs, deriving $v_{t}$.
-Figure 3: Airfoil generation by FuncGenFoil.
-We first sample latent function from a zero-mean Gaussian process with kernel function $K$ at any resolution.
-Then we solve an ODE flow through $t=0$ to $t=1$ using numerical ODE solver using velocity operator $v_{\theta}$ along this generation process and can generate diverse airfoils.
-As shown in (c), A single FuncGenFoil model can generate airfoils at arbitrary resolutions, from low to high, while guaranteeing identical outputs when sampling from the same latent function—an inherent property of the model’s architecture.
-Airfoil Geometry Parametrization.
-Since an airfoil has circular topology, we introduce the variable $\alpha\sim[0,2\pi]$ for convenience.
-We denote $y(\alpha)=u_{1}(\alpha)$ and $x(\alpha)=\frac{\cos(2\pi\alpha)+1}{2}$ [^1] For the remainder of this paper, we omit the explicit dependence on $\alpha$ for clarity., achieving the function of airfoil $u_{0}=(x(\alpha),y(\alpha))$.
-Training. We train $v_{\theta}$ under the denoising training process. Given an airfoil geometry $u_{1}$, we compute its corresponding noised sample at time $t$ as follows:
-$u_{t}=t\times u_{1}+(1-t)\times u_{0}$ (2)
-Then we feed $u_{t}$ into $v_{\theta}$ and train $v_{\theta}$ by matching the velocity operator:
-$\mathcal{L}_{\mathrm{FM}}=E_{t\sim[0,1],u_{0},u_{1}}\left[\|v_{t}-v_{\theta}(u_{t},c,t)\|^{2}\right],$ (3)
-where $v_{t}$ is the ground-truth velocity:
-$v_{t}=u_{1}-u_{0}.$ (4)
-Detailed algorithm for training velocity operator is shown in Appendix B Algorithm 2.
+The velocity operator $v_{t}$ gradually transforms a latent function $u_{0}$ sampled from a stochastic process $\mathcal{P}$ into an airfoil function $u_{1}$ belonging to the target airfoil distribution $\mathcal{Q}$.
+The velocity operator is a key component, as it must handle function inputs consisting of point sets with arbitrary resolution or positions.
+Specifically, the operator takes as input the partially generated airfoil $u_{t}=[u_{t}(\alpha_{0}),u_{t}(\alpha_{1}),u_{t}(\alpha_{2}),\dots]$ and outputs a velocity function corresponding to these positions, $[v_{t}(\alpha_{0}),v_{t}(\alpha_{1}),v_{t}(\alpha_{2}),\dots]$.
+We realize this by establishing a parameterized neural operator $v_{\theta}$ with model weights $\theta$, constructed as a Neural Operator model capable of processing function-space data at arbitrary resolutions in a single unified model, as shown in Figure 2.
+Specifically, FuncGenFoil acts as a conditional continuous-time generative model, where the velocity operator $v_{\theta}(u_{t},c,t)$ takes the noised airfoil $u_{t}$, optional conditioning variables $c$, and generation timestamp $t$ as inputs, and consistently outputs the deformation $v_{t}$ as a function.
+We train $v_{\theta}(u_{t},c,t)$ using Operator Flow Matching [53].
+Figure 2: Top: Overview of FuncGenFoil’s neural network and training scheme.
+The model is a Fourier Neural Operator designed for point cloud data, although any other neural operator capable of general-purpose function-space approximation may be used.
+The model takes as input a function $u_{t}$ (point cloud data at an arbitrary resolution $d$), optional design condition variables $c$, and the generation time $t$.
+It then processes this input function consistently and outputs the current velocity operator $v_{\theta}(u_{t},c,t)$ for calculating the flow matching loss.
+Bottom: Inference with FuncGenFoil is conducted by first sampling a random latent function from a Gaussian process and then reconstructing the airfoil by solving an ODE.
+Training. We train $v_{\theta}$ under the simplest denoising training process. Given an airfoil geometry $u_{1}$, we compute its corresponding noised sample at time $t$ as follows:
+$u_{t}|_{u_{0},u_{1}}=t\times u_{1}+(1-t)\times u_{0}$ (1)
+and then the ground-truth velocity $v_{t}(u_{t})|_{u_{0},u_{1}}$ given $u_{0}$ and $u_{1}$ can be computed as:
+$v_{t}(u_{t})|_{u_{0},u_{1}}=\frac{du_{t}}{dt}=u_{1}-u_{0}.$ (2)
+We can train $v_{\theta}$ by matching the velocity operator using Flow Matching loss:
+$\mathcal{L}_{\mathrm{FM}}=E_{t\sim[0,1],u_{t}}\left[\|v_{t}(u_{t})-v_{\theta}(u_{t},c,t)\|^{2}\right],$ (3)
+which according to Conditional Matching Theorem, has same gradients with Conditional Flow Matching loss:
+$\mathcal{L}_{\mathrm{CFM}}=E_{t\sim[0,1],u_{0},u_{1}}\left[\|v_{t}(u_{t})|_{u_{0},u_{1}}-v_{\theta}(u_{t}|_{u_{0},u_{1}},c,t)\|^{2}\right],$ (4)
+$\nabla_{\theta}\mathcal{L}_{\mathrm{FM}}(\theta)=\nabla_{\theta}\mathcal{L}_{\mathrm{CFM}}(\theta).$ (5)
+See more proofs in Chapter 4 of Flow Matching Guide and Code [36].
 Inference.
 Given a trained velocity operator $v_{\theta}$, the inference process, i.e., airfoil generation process, is equal to deriving airfoil geometry at time $t=1$, denoted as $u_{1}$, based on a latent coding $u_{0}$ sampled from the stochastic process $\mathcal{P}$.
-$\mathcal{P}$ is assumed as a Gaussian Process $\mathcal{GP}(0,K)$, where $K$ is a covariance kernel function.
-Therefore, $u_{1}$ could be derived by solving the ODE in Equation 1 numerically (e.g., using the Euler method) as follows:
+$\mathcal{P}$ is assumed as a Gaussian Process $\mathcal{GP}(0,K)$ in this work, where $K$ is a covariance kernel function.
+Therefore, $u_{1}$ could be derived by solving the generation ODE numerically (e.g., using the Euler method) as follows:
 $u_{1}=u_{0}+\int_{t=0}^{t=1}v_{\theta}(u_{t},c,t)\,\mathrm{d}t.$ (6)
-The generation process is shown in Fig. 3 while the detailed algorithm is shown in Appendix B Algorithm 3.
+Detailed implementations for training velocity operator and model inference are shown in Appendix B, Algorithm 1 and Algorithm 2.
 3.2 Airfoil Editing
@@
 After a few iterations of fine-tuning, we can generate edited airfoil $u_{1}^{\delta}$ with high accuracy.
-The airfoil editing task enables the user to modify parts of the geometry of a given airfoil, effectively generating a new airfoil geometry while preserving the user-edited sections, denoted as $u_{1}^{\delta}$.
-We achieve this by an optimization method that maximum a posteriori (MAP) estimation, $\max p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ where $p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ is a probabilistic model that constrains the optimized airfoil $u_{1}^{{}^{\prime}}$ fulfilling the editing requirements and following the generation prior.
+The airfoil editing task enables the user to modify parts of the geometry of a given airfoil $u_{1}$, effectively generating a new airfoil geometry $u_{1}^{{}^{\prime}}$ while preserving the user-edited sections, denoted as $u_{1}^{\delta}$.
+We achieve this by finetuning the pretrained model via maximum a posteriori (MAP) estimation, $\max p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ where $p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})$ is a probabilistic model that constrains the optimized airfoil $u_{1}^{{}^{\prime}}$ fulfilling the editing requirements and following the generation prior.
 To achieve the constraint probability model, we disentangled it with Bayes’ Rule:
-$\max_{u_{1}^{{}^{\prime}}}p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})\;=\;\frac{p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})\cdot p(u_{1}^{{}^{\prime}})}{p(u_{1}^{\delta})}$ (6)
-$\Rightarrow$ $\max_{u_{1}^{{}^{\prime}}}\log{p(u_{1}^{\delta}|u_{1}^{{}^{\prime}})}+\log{p(u_{1}^{{}^{\prime}})}-\log{p(u_{1}^{\delta})},$
+$\max_{u_{1}^{{}^{\prime}}}p(u_{1}^{{}^{\prime}}\mid u_{1}^{\delta})\;=\;\frac{p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})\cdot p(u_{1}^{{}^{\prime}})}{p(u_{1}^{\delta})}\Rightarrow\max_{u_{1}^{{}^{\prime}}}\log{p(u_{1}^{\delta}|u_{1}^{{}^{\prime}})}+\log{p(u_{1}^{{}^{\prime}})}-\log{p(u_{1}^{\delta})},$ (7)
 where $p(u_{1}^{\delta}\mid u_{1}^{{}^{\prime}})$ is a Gaussian measure, so its log term becomes a Mean Square Error (MSE) between $u_{1}^{\delta}$ and $u_{1}^{{}^{\prime}}$, constraining the user edited parts in $u_{1}^{{}^{\prime}}$ keeping consistent with $u_{1}^{\delta}$.
-$p(u_{1}^{{}^{\prime}})$ is the prior supported by the trained generative model so its log value is derived by Hutchinson trace estimator Hutchinson (1990); Grathwohl et al.
-(2019) effectively.
+We can sample $u_{1}^{{}^{\prime}}$ through solving the velocity operator using Neural ODE method[6] and calculate $\log p(u_{1}^{{}^{\prime}})$ at the same time using FFJORD method [17] with Hutchinson trace estimator [22].
+$p(u_{1}^{{}^{\prime}})$ is the prior supported by the trained generative model.
 $p(u_{1}^{\delta})$ is the marginal likelihood, which does not depend on $u_{1}^{{}^{\prime}}$.
@@
 $\max_{u_{1}^{{}^{\prime}}}\frac{1}{2\sigma^{2}}\sum_{i\in\Delta}(u_{1}^{\delta,i}-u_{1}^{{}^{\prime},i})^{2}+\log{p(u_{1}^{{}^{\prime}})},$ (8)
-where $\sigma$ is noise, $\Delta$ denotes point indices in edited part $u_{1}^{\delta}$.
+where $\sigma$ is noise scale for editing, $\Delta$ denotes point indices in edited part $u_{1}^{\delta}$.
 For more realistic generation results, the optimization does not directly adjust $u_{1}^{{}^{\prime}}$; instead, we optimize $u_{1}^{{}^{\prime}}$ indirectly by fine-tuning the entire generative model for a few iterations.
-The fine-tuning process is illustrated in Fig.
-4, while its details are provided in Alg.
-1 in the main text.
+The fine-tuning process is illustrated in Figure 3, while its details are provided in Alg.
+3 in the Appendix.
 Specifically, we first initialize $u^{\prime}_{1}$ as a resample data of $u_{1}$, by extracting its latent code, denoted as $\alpha_{\theta}$, via the inverse of our generative model and re-generating $u_{1}^{{}^{\prime}}$ from this code.
@@
 After the model training, the new $u_{1}^{{}^{\prime}}$ is the new edited generation results.
-Algorithm 1 Airfoil Editing.
-Input: pretrained neural operator $v_{\theta}$, original airfoil function $u_{1}$ (optional) or latent function $u_{0}$ (optional), editing requirement $\delta$, editing resolution $d$, sampling time steps $T$ and steps length $\mathrm{d}t$.
-Parameter: gaussian process $\mathcal{GP}(0,K)$ for sampling $u_{0}$, noise level $\sigma$.
-Output: new airfoil $\{y_{i}\}$ at resolution $d$.
-1:  Inversely sample original airfoil function $\{u_{1,i}\}$ through $v_{\theta}$ and get its latent function $\{u_{0,i}\}$.
-2:  Set $\{a_{\theta,i}\}=\{u_{0,i}\}$.
-3:  while finetuning… do
-4:     Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
-5:     Compute $\mathcal{L}_{\mathrm{MAP}}$.
-6:     Compute gradient and update $\theta$.
-7:  end while
-8:  Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
-9:  return $\{y_{i}\}=\{u_{1,i}^{{}^{\prime}}\}$
 4 Experiments
-4.1 Experiment Settings
+4.1 Experimental Settings
 Tasks.
-We evaluate two tasks in the airfoil inverse design problem: conditional airfoil generation and freestyle airfoil editing.
-In the conditional generation task, the model is given a set of 11 geometric parameters that describe the airfoil geometry.
-Detailed parameter definitions are provided in the Supplementary Materials.
+We evaluate two tasks within the airfoil inverse design problem: conditional airfoil generation and freestyle airfoil editing.
+In the conditional generation task, the model is provided with a set of 11 geometric parameters describing the airfoil geometry.
+Detailed parameter definitions are provided in Table 6 in the Appendix.
 The model must generate airfoils that satisfy these geometric constraints.
-In the freestyle editing task, the model takes an original airfoil and a target modification, such as adjusting the position of a specific point on the airfoil.
-The selected point can be any location on the airfoil surface.
-The model must generate an airfoil that reflects the specified modification.
-Metrics. We adopt the metrics introduced in AFBench Liu et al. (2024) to evaluate the generated airfoils:
-Label Error measures the difference between the physical parameters of the generated or edited airfoil and the intended target parameters, calculated as $\sigma_{i}=\left|\hat{p}_{i}-p_{i}\right|$, where $\sigma_{i}$ is the label error for the $i$-th parameter, $\hat{p}_{i}$ is the $i$-th geometric parameter from the generated airfoil, and $p_{i}$ is the corresponding target parameter.
-Smaller values indicate better alignment with the target parameters.
-Diversity quantifies the variety of generated airfoils, calculated as $D=\frac{1}{n}\sum_{i=1}^{n}\log\det(L_{S_{i}})$, where $n$ is the number of samples, and $L_{S_{i}}$ is the similarity matrix of the $i$-th subset, computed based on Euclidean distances between airfoils in the subset.
+In the freestyle editing task, the model receives an original airfoil and a target modification, such as adjusting the position of a specific point on the airfoil.
+The selected point can be located anywhere on the airfoil surface.
+The model must generate an airfoil reflecting the specified modification.
+Metrics. We adopt the metrics introduced in AFBench [37] to evaluate the generated airfoils:
+Label Error measures the difference between the PARSEC parameters of the generated or edited airfoil and the intended target parameters, calculated as $\sigma_{i}=\left|\hat{p}_{i}-p_{i}\right|$, where $\sigma_{i}$ is the label error for the $i$-th parameter, $\hat{p}_{i}$ is the $i$-th geometric parameter of the generated airfoil, and $p_{i}$ is the corresponding target parameter.
+Smaller values indicate better alignment with target parameters.
+To summarize all $11$ label errors, we present both the arithmetic mean $\bar{\sigma}_{a}$ and the geometric mean $\bar{\sigma}_{g}$ for absolute and relative average errors.
+Diversity quantifies the variety among generated airfoils, calculated as $D=\frac{1}{n}\sum_{i=1}^{n}\log\det(L_{S_{i}})$, where $n$ is the number of subsets, and $L_{S_{i}}$ is the similarity matrix of the $i$-th subset, computed based on Euclidean distances between airfoils within the subset.
 Higher values indicate greater diversity among generated airfoils.
-Smoothness measures the geometric smoothness of the generated airfoils, calculated as:
-$\small M=\sum_{i=1}^{N}\text{Distance}(P_{n},P_{n-1}P_{n+1}),$ (8)
-where $P_{n}$ is the $n$-th point, and $P_{n-1}P_{n+1}$ is the line segment between adjacent points.
+Smoothness measures the geometric smoothness of generated airfoils, calculated as:
+$\small M=\sum_{n=1}^{N}\text{Distance}(P_{n},P_{n-1}P_{n+1}),$ (9)
+where $P_{n}$ is the $n$-th point, and $P_{n-1}P_{n+1}$ is the line segment connecting its adjacent points.
 The function $\text{Distance}(P_{n},P_{n-1}P_{n+1})$ computes the perpendicular distance from $P_{n}$ to this line segment.
@@
 Datasets.
-To benchmark our method, we conduct experiments on three datasets: UIUC Selig (1996), Supercritical Airfoil (Super), and AF-200K.
+To benchmark our method, we conduct experiments on three datasets: UIUC [50], Supercritical Airfoil (Super), and AF-200K.
 UIUC contains 1,600 designed airfoil geometries.
@@
 Baselines.
-We include the baseline models proposed in AFBench Liu et al.
-(2024), specifically the conditional VAE (CVAE), conditional GAN (CGAN), the modified VAE with PARSEC parameters and control keypoints (PK-VAE), along with PK-GAN, PKVAE-GAN, the U-Net-based PK-DIFF, and the transformer-based PK-DIT.
-Table 1: Quantitative comparison between FuncGenFoil and baseline methods across different datasets for conditional generation task.
-Label error, diversity, and smoothness of the generated airfoils are reported.
+We compare FuncGenFoil with baseline models proposed in AFBench [37], specifically the conditional VAE (CVAE), conditional GAN (CGAN), the modified VAE with PARSEC parameters and control keypoints (PK-VAE), as well as PK-GAN, PKVAE-GAN, the U-Net-based PK-DIFF, and the transformer-based PK-DIT.
+Table 1: Quantitative comparison between FuncGenFoil and baseline methods across different datasets for the conditional generation task.
+Label error, diversity, and smoothness of generated airfoils are reported.
 […]
-$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
-CVAE AF-200K 72.9 52.5 35.2 15900 99 95 29000 19.1 15.3 46 104 4131 -155.4 7.09
-CGAN AF-200K 107 85.0 54.4 23200 143 137 59600 25.3 22.3 53 129 7596 -120.5 7.31
-PK-VAE AF-200K 63.0 47.9 31.3 8620 66 64 17100 13.5 9.3 33 78 2375 -150.1 5.93
-PK-GAN AF-200K 81.8 63.0 47.0 21030 120 117 32470 22.5 19.6 50 122 4923 -112.3 3.98
-PKVAE-GAN AF-200K 56.8 31.7 31.0 5650 46 43 12000 9.1 5.1 28 63 1633 -129.6 2.89
-PK-DIT AF-200K 11.2 32.3 15.4 1050 13 11.5 9790 0.5 0.5 23 24 997 -93.2 1.04
-FuncGenFoil AF-200K 1.84 17.4 0.56 721 47.7 0.98 1676 0.45 0.65 160 174 255 -71.6 1.41
-PK-DIT UIUC 63.8 51.4 33.6 11830 87 84.9 25700 16.9 11.9 36 98 3456 -141.7 6.03
-FuncGenFoil UIUC 14.2 12.7 0.86 434.9 90.7 3.7 3559 0.55 0.69 129.9 135.2 398.4 -66.3 1.17
-PK-DIT Super 52.0 35.0 24.0 3010 29 33.2 10500 8.3 2.6 27 33 1250 -123.4 1.97
-FuncGenFoil Super 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
-Table 2: Quantitative evaluation of the airfoil editing task across different edit scales.
-The mean squared error (MSE) between the generated and target airfoils as well as the smoothness of the generated airfoils are reported.
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+CVAE AF-200K 72.9 52.5 35.2 15900 99 95 29000 19.1 15.3 46 104 4131 149.2 -155.4 7.09
+CGAN AF-200K 107 85.0 54.4 23200 143 137 59600 25.3 22.3 53 129 7596 217.6 -120.5 7.31
+PK-VAE AF-200K 63.0 47.9 31.3 8620 66 64 17100 13.5 9.3 33 78 2375 106.0 -150.1 5.93
+PK-GAN AF-200K 81.8 63.0 47.0 21030 120 117 32470 22.5 19.6 50 122 4922 179.5 -112.3 3.98
+PKVAE-GAN AF-200K 56.8 31.7 31.0 5650 46 43 12000 9.1 5.1 28 63 1633 77.6 -129.6 2.89
+PK-DIT AF-200K 11.2 32.3 15.4 1050 13 11.5 9790 0.5 0.5 23 24 997 23.5 -93.2 1.04
+FuncGenFoil AF-200K 1.84 17.4 0.56 721 47.7 0.98 1676 0.45 0.65 160 174 255 14.9 -71.6 1.41
+PK-DIT UIUC 63.8 51.4 33.6 11830 87 84.9 25700 16.9 11.9 36 98 3456 129.8 -141.7 6.03
+FuncGenFoil UIUC 11.6 11.8 0.38 698 23.5 0.56 12339 0.27 0.44 110.5 115.0 1210 15.1 -91.1 1.33
+PK-DIT Super 52.0 35.0 24.0 3010 29 33.2 10500 8.3 2.6 27 33 1250 58.1 -123.4 1.97
+FuncGenFoil Super 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
+Table 2: Quantitative evaluation of the FuncGenFoil model across different sampling super-resolutions for the conditional generation task. The training resolution is $257$.
+[…]
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+Super 257 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
+513 0.71 8.19 0.12 200.8 4.81 0.12 174.7 0.09 0.15 48.2 47.8 44.1 3.26 -97.5 0.52
+1025 0.71 8.29 0.12 202.9 4.93 0.12 175.4 0.09 0.15 61.4 57.8 46.5 3.44 -91.1 0.37
+Implementation Details.
+On the AF-200K dataset, we trained for 2 million iterations with a batch size of 2,048 using 8 NVIDIA 4090 GPUs.
+On the Supercritical Airfoil and the UIUC dataset, we trained for 1 million iterations with a batch size of 1,024 on 4 single NVIDIA 4090 GPUs.
+We use Gaussian processes with a Matérn kernel function ($\nu=2.5,l=0.03$) as the prior.
+Other training hyperparameters are detailed in Appendix C.
+4.2 Main Results
+Conditional Airfoil Generation.
+As shown in Table 1, FuncGenFoil outperforms the strongest baseline method (PK-DIT) across all metrics.
+For label error, FuncGenFoil achieves reductions in arithmetic mean label error of 74.4% on AF-200K, 65.0% on UIUC, and 96.6% on Super.
+More importantly, it also reduces the geometric mean label error by 36.6% on AF-200K, 88.4% on UIUC, and 94.7% on Super.
+This substantial decrease in label error underscores the effectiveness of our approach for generating airfoils that more precisely adhere to target geometric parameters.
+In terms of diversity, FuncGenFoil demonstrates notable improvements, surpassing the best baseline methods by 23.2%, 35.7%, and 15.8% on AF-200K, UIUC, and Super, respectively.
+This highlights the model’s superior capability in capturing and generating a broader spectrum of valid airfoil designs.
+Additionally, the generated airfoils exhibit enhanced surface smoothness, as evidenced by reductions of 4.70 and 0.96 in smoothness values ($10^{-2}$) on the UIUC and Super datasets, respectively.
+This improvement is particularly crucial for aerodynamic performance, as smoother airfoil surfaces contribute to reduced drag and improved flow characteristics.
+Table 3: Quantitative evaluation of the airfoil editing task across different editing scales.
+Mean squared error (MSE) between generated and target airfoils and smoothness of generated airfoils are reported.
 […]
@@
 0.0004 2.75 1.15
-Table 3: Quantitative evaluation of our model (training resolution=257) across different sampling resolutions for conditional generation task.
-[…]
-$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}$
-Super 257 0.73 1.9 0.09 53.2 1.07 0.07 56.4 0.06 0.06 10.5 13.6 12.5 -108.8 0.98
-513 0.74 2.59 0.13 55.1 1.54 0.09 57.7 0.07 0.08 41.1 40.0 18.1 -100.5 0.49
-1025 0.75 3.18 0.17 56.5 2.03 0.12 59.2 0.07 0.09 60.8 55.0 21.6 -92.9 0.31
-Implementation Details.
-On the AF-200K dataset, we trained for 2 million iterations with a batch size of 2,048 using 8 NVIDIA 4090 GPUs.
-On the Supercritical Airfoil dataset, we trained for 1 million iterations with a batch size of 256 on a single NVIDIA 4090 GPU.
-We use the Matérn kernel function as the kernel for the Gaussian process at time $t=0$.
-Other training hyperparameters are detailed in the Supplementary Materials.
-4.2 Main Results
-Conditional Airfoil Generation.
-As shown in Table 1, FuncGenFoil significantly outperforms the strongest baseline method (PK-DIT) across all metrics.
-For label error, FuncGenFoil achieves a reduction in average error by 88.5% on UIUC, 99% on Super, and 74.4% on AF-200K.
-This substantial decrease in label error underscores the effectiveness of our approach in generating airfoils that more precisely adhere to target geometric parameters.
-In terms of diversity, FuncGenFoil demonstrates notable improvements, surpassing the best baseline methods by 75.4, 14.6, and 21.6 on UIUC, Super, and AF-200K, respectively.
-This highlights the model’s superior capability in capturing and generating a broader spectrum of valid airfoil designs.
-Additionally, our generated airfoils exhibit enhanced surface smoothness, as evidenced by reductions of 0.99 and 4.86 in smoothness values ($10^{-2}$) on the UIUC and Super datasets, respectively.
-This improvement is particularly crucial for aerodynamic performance, as smoother airfoil surfaces contribute to reduced drag and improved flow characteristics.
 Freestyle Airfoil Editing.
-Table 2 shows the average performance across 300 editing cases using FuncGenFoil.
-In each case, the model adjusts 2 to 4 randomly selected positions on the wing surface to target locations over 10 fine-tuning steps, with edit scales ranging from $1\times 10^{-4}$ to $4\times 10^{-4}$.
-The results demonstrate that FuncGenFoil can perform accurate airfoil editing with minimal errors (less than $2.75\times 10^{-7}$ MSE) and high surface smoothness (less than 1.16 ×$\times 10^{-2}$ smoothness value) with just a few fine-tuning steps.
-Fig.
-5 illustrates the editing requirements and generated airfoils during the fine-tuning stage.
-It shows that our model faithfully completes the freestyle editing task by generating reasonable airfoils according to user-specified editing requirements.
-Figure 5: Examples of two FuncGenFoil instances performing airfoil editing over 100 training iterations.
-The red section represents the editing requirement, while the black airfoil shows the generated samples at each training stage.
+Table 3 shows average performance over 300 editing cases using FuncGenFoil.
+In each case, the model adjusts 2 to 4 randomly selected positions on the airfoil surface to target locations over 10 fine-tuning steps, with editing scales ranging from $1\times 10^{-4}$ to $4\times 10^{-4}$.
+Results demonstrate that FuncGenFoil achieves accurate airfoil editing with minimal errors (less than $2.75\times 10^{-7}$ MSE) and high surface smoothness (less than 1.16 ×$\times 10^{-2}$ smoothness value) after only a few fine-tuning steps.
+Figure 4 illustrates example editing requirements and corresponding generated airfoils during the fine-tuning stage, demonstrating that our model effectively completes freestyle editing tasks by generating accurate airfoils according to user-specified edits.
+Figure 4: Examples of eight different FuncGenFoil instances performing airfoil editing over 20 training iterations.
+The orange points and sections represent the editing requirements as constraints; the gray region is the original airfoil, while the blue region shows the generated edited airfoil.
 The results demonstrate that the generated airfoil quickly adapts to the editing requirements within a few iterations, achieving a natural and smooth function regression.
+The editing scheme can be completely customized according to the user’s preference.
 Any-Resolution Airfoil Generation.
@@
 The results are presented in Table 2.
-We observe consistent generation quality at $\times$ and even $\times$ the training resolution, with a maximum increase of 9.1×$\times 10^{-3}$ in average label error, a decrease of 16.1 in diversity, and an increase of 0.67×$\times 10^{-2}$ in smoothness value.
+We observe consistent generation quality at $\times$ and even $\times$ the training resolution, with a maximum increase of 4.6×$\times 10^{-3}$ in average label error, a decrease of 12.8 in diversity, and an increase of 0.64×$\times 10^{-2}$ in smoothness value.
+Figure 5: Left: Comparison of the lift-to-drag ratio histograms for CRM wings in the original dataset and the samples generated by our FuncGenFoil model.
+Right: Visualization of the pressure coefficient for the generated CRM wings, obtained through aerodynamic simulation.
 Aerodynamic Simulation.
 To further assess the physical properties of the generated airfoils and validate the effectiveness of our method, we analyze their aerodynamic performance using the NASA Common Research Model (CRM)  [^2] https://commonresearchmodel.larc.nasa.gov/ and perform Reynolds-Averaged Navier-Stokes (RANS) computational fluid dynamics (CFD) simulations on the generated samples.
-The CRM dataset contains 135,000 CRM wing geometries along with their corresponding aerodynamic performance, computed using the RANS CFD solver ADflow Mader et al.
-(2020).
+The CRM dataset contains 135,000 CRM wing geometries along with their corresponding aerodynamic performance, computed using the RANS CFD solver ADflow [39].
 We pretrain the FuncGenFoil model on all 135,000 CRM wing geometries and generate 500 new CRM wing geometries for RANS CFD evaluation.
-We analyze the lift-to-drag ratio ($L/D$) of these newly generated geometries and compare them with the original dataset, as shown in Fig.
-6.
+We analyze the lift-to-drag ratio ($L/D$) of these newly generated geometries and compare them with the original dataset, as shown in Figure 5.
 Our results indicate that the $L/D$ distribution of the generated samples closely aligns with that of the original dataset, with the highest density occurring around $L/D=19$, demonstrating the model’s ability to learn and generate physically plausible wing geometries.
-Additionally, we visualize the coefficient of pressure contours for selected CFD cases in Fig.
-7.
+Additionally, we visualize the coefficient of pressure contours for selected CFD cases in Figure 5.
 These visualizations confirm that FuncGenFoil can generate airfoils with diverse aerodynamic performance characteristics.
-Figure 6: Comparison of the lift-to-drag ratio histograms for CRM wings in the original dataset and the samples generated by our FuncGenFoil model.
-Figure 7: Visualization of the pressure coefficient for the generated CRM wings, obtained through aerodynamic simulation.
 4.3 Ablation Study
-We first ablate the use of the Matérn kernel $K$ with different parameters as the kernel function for the sampled Gaussian process.
-As shown in Table 4, we compare the cases of Matérn kernel $\nu=1.5$, $l=0.01$, and Matérn kernel $\nu=2.5$, $l=0.03$, where the latter is smoother due to the combined effect of both parameters.
-Notably, the limits of $l=0$ and $\nu\to\infty$ correspond to pure white noise and the RBF kernel, respectively.
-To further highlight the importance of kernel function selection, we also test two models of different sizes (by setting different FNO modes).
-From Table 4, we observe that a smoother kernel improves both the smoothness and accuracy of the generated airfoils.
-Additionally, the comparison highlights that, under our setup, selecting a smoother kernel leads to a more substantial performance improvement than increasing the model size.
-This underscores the benefit of modeling the function object in the airfoil generation task rather than modeling discrete points.
-In Table 5, we compare the MSE of constraint condition calculations in the airfoil editing task using two initialization schemes: (1) using the latent variables obtained through ODE inversion of the original airfoil as the prior, and (2) using a zero prior $u_{0}=0$ without ODE inversion.
-Both schemes are fine-tuned with the same number of iterations.
-As shown, the ODE inversion-based prior significantly reduces the MSE compared to the zero prior, with a maximum reduction of 84.85.
-This underscores the importance of incorporating original airfoil information during the initialization process.
-Table 4: Ablation study on the impact of different kernels and model sizes for the conditional airfoil generation task.
-Matérn $\nu$ Matérn $l$ FNO modes $\bar{\sigma}\downarrow$ $\mathcal{M}\downarrow$
-1.5 0.01 32 0.108 0.0119
-2.5 0.03 32 0.013 0.0098
-1.5 0.01 48 0.039 0.0101
-2.5 0.03 48 0.013 0.0098
+Figure 6: (A): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using kernels of different smoothness but same characteristic length $l=0.03$.
+(B): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using Matérn kernel $\nu=3.5$ with different characteristic length $l$.
+(C): Comparison of the geometric mean error $\bar{\sigma}_{g}$ using different Fourier mode in FNO using Matérn kernel of same $l=0.03$.
+Kernels of Gaussian Process.
+The type of Gaussian process kernel used as a prior in the model influences the function space into which the data is diffused.
+We perform ablation experiments across several kernel types, including the white noise kernel, the Matérn kernel with smoothness parameter $\nu\in[1.5,2.5,3.5]$, and the radial basis function (RBF) kernel.
+We measure the geometric mean $\bar{\sigma}_{g}$ under different smoothness, characteristic length, and Fourier mode settings, as shown in Figure 6.
+We observe that FunGenfoil benefits from employing smoother kernels, which enhance both the smoothness and accuracy of the generated airfoils.
+This effect arises because the parameter $\nu$ primarily controls the smoothness of the function space, and the RBF kernel corresponds to the case $\nu\rightarrow\infty$.
+Additionally, we find that there exists an optimal characteristic length at $l=0.03$ for the kernel.
+In our experiments, FunGenfoil does not exhibit improved performance with an increased number of Fourier modes.
+This phenomenon may be due to the predominantly low-frequency characteristics of supercritical airfoils, such that adding higher modes in the FNO layer introduces unnecessary model complexity.
+More detailed experimental results are provided in Table 8 in the Appendix.
 Table 4: Ablation study on the effect of different latent variable initialization methods for the airfoil editing task.
 Edit scale MSE$\mathcal{\downarrow}(10^{-7})$
-w/ ODE inversion w/o ODE inversion
+w/ ODE inv. w/o ODE inv.
 0.0001 2.42 ($\downarrow$83.88) 86.3
@@
 0.0004 2.75 ($\downarrow$84.85) 87.6
+Editing with Prior via ODE Inversion.
+In Table 4, we compare MSE of constraint condition calculations in the airfoil editing task using two initialization schemes: (1) using the latent variables obtained through ODE inversion of the original airfoil as the prior, and (2) using a zero prior $u_{0}=0$ without ODE inversion.
+Both schemes are fine-tuned with the same number of iterations.
+The ODE inversion-based prior significantly reduces the MSE compared to the zero prior, with a maximum reduction of 84.85.
+This underscores the importance of incorporating original airfoil information during the initialization process.
+ODE Numerical Method.
+Model inference involves solving an ordinary differential equation (ODE), and the choice of numerical integration scheme affects its performance.
+Using the same velocity operator $v_{\theta}$ trained on the supercritical airfoil dataset and sampling from a Gaussian prior with a Matérn kernel ($\nu=2.5$, $l=0.03$), we evaluate FuncGenFoil across various time step sizes and ODE solvers (Euler, midpoint, and fourth-order Runge-Kutta (RK4) methods).
+Results in Table 5 show that FuncGenFoil benefits from larger time steps using Euler method.
+However, RK4 may degrade performance at excessively small step sizes, potentially because the model enters regions where the velocity operator is inadequately trained.
+Table 5: Quantitative experiments on the impact of ODE solver and time step on sampling quality.
+[…]
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+Euler $10$ 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
+$50$ 0.62 3.61 0.07 92.5 2.08 0.06 96.8 0.06 0.07 14.5 15.7 20.6 1.60 -107.6 0.99
+$100$ 0.67 3.13 0.09 83.9 1.84 0.08 90.6 0.07 0.06 12.3 13.1 18.7 1.55 -106.4 0.99
+Midpoint $10$ 0.68 5.29 0.08 161.3 3.12 0.08 178.9 0.06 0.09 21.4 22.4 35.8 2.17 -105.1 1.00
+$50$ 0.75 3.82 0.12 137.9 2.34 0.10 153.4 0.08 0.06 14.1 14.9 29.8 1.97 -103.3 0.99
+$100$ 0.82 3.94 0.14 125.7 2.22 0.13 137.4 0.08 0.07 14.4 14.2 27.2 2.06 -101.6 0.99
+RK4 $10$ 15.9 46.3 1.21 4406 25.7 1.24 4795 1.17 1.68 333 396 911 36.6 -67.5 3.05
+$50$ 0.92 7.63 0.24 482.7 5.55 0.22 511.8 0.16 0.19 39.3 41.3 99.1 4.68 -92.9 1.07
+$100$ 0.82 5.65 0.18 160.6 3.21 0.17 170.8 0.10 0.12 22.0 22.8 35.1 2.78 -96.9 1.00
 5 Conclusion
@@
 This work paves the way for more efficient, scalable, and versatile airfoil generation, with significant applications in aerodynamic shape optimization for aircraft manufacturing.
-Appendix A Airfoil Geometry Parameters
-Table 6: Geometric parameters of the airfoils used in conditional generation tasks
-Index Symbol Meaning
-1 $R_{\mathrm{le}}$ the leading edge radius
-2 $X_{\mathrm{up}}$ upper crest position x
-3 $Y_{\mathrm{up}}$ upper crest position y
-4 $Z_{\mathrm{xxup}}$ upper crest curvature
-5 $X_{\mathrm{lo}}$ lower crest position x
-6 $Y_{\mathrm{lo}}$ lower crest position y
-7 $Z_{\mathrm{xxlo}}$ lower crest curvature
-8 $Y_{\mathrm{te}}$ trailing edge position
-9 $\Delta Y_{\mathrm{te}}$ trailing thickness
+Appendix A Problem Setting for Airfoil Generation
+A.1 Airfoil Parametrization
+In this paper, we perform a circular parameterization of the airfoil curves, as illustrated in Figure 7.
+Specifically, we project a unit circle onto the $x$-axis and adopt the angle $\alpha$ as the function domain.
+The parameterized airfoil function begins at the trailing edge on the suction side, where $\alpha=0$, and ends at the trailing edge on the pressure side, where $\alpha=1$.
+In this manner, we transform the airfoil data from a circular topology into a well-defined functional form suitable for reconstruction by generative models.
+Figure 7: Illustration of the airfoil curve parameterization employed in FuncGenFoil.
+A.2 Design Parameters of Airfoil Geometry
+Index Symbol Geometric Meaning
+1 $\mathrm{R_{le}}$ leading edge radius
+2 $\mathrm{X_{up}}$ upper crest position x
+3 $\mathrm{Y_{up}}$ upper crest position y
+4 $\mathrm{Z_{xxup}}$ upper crest curvature
+5 $\mathrm{X_{lo}}$ lower crest position x
+6 $\mathrm{Y_{lo}}$ lower crest position y
+7 $\mathrm{Z_{xxlo}}$ lower crest curvature
+8 $\mathrm{Y_{te}}$ trailing edge position
+9 $\Delta\mathrm{Y_{te}}$ trailing thickness
 10 $\alpha_{\mathrm{te}}$ trailing edge angle up
 11 $\beta_{\mathrm{te}}$ trailing edge angle down
-Algorithm 2 Generative Model Training.
+(a) PARSEC parameters and its geometric meaning.
+(b) Demonstration of PARSEC parameters.
+Table 6: Eleven PARSEC parameterization in airfoil design.
+Algorithm 1 Model Training.
 Input: data resolution $d$, data $u_{1}$, design condition variables $c$ (optional).
@@
 9:  return $v_{\theta}$.
-Algorithm 3 Inference Process (Airfoil Generation)
+Algorithm 2 Model Inference.
 Input: sampling resolution $d$, sampling time steps $T$ and steps length $\mathrm{d}t$, latent function $u_{0}$ (optional), design condition variables $c$ (optional).
@@
 1:  Let $t=0$, $u_{0}\sim\mathcal{GP}(0,K)$ at resolution $d$ and get $\{u_{0,i}\}$.
-2:  while $t<=1$ do
+2:  while $t\leq 1$ do
 3:     Compute $v_{\theta}(u_{t},c,t)$ at resolution $d$ and get $\{v_{\theta,i}\}$.
@@
 7:  return $\{y_{i}\}=\{u_{1,i}\}$
-Table 7: Training hyperparameters and model parameters in conditional airfoils generation tasks
+Algorithm 3 Model Finetuning (Airfoil Editing).
+Input: pretrained neural operator $v_{\theta}$, original airfoil function $u_{1}$ (optional) or latent function $u_{0}$ (optional), editing requirement $\delta$, editing resolution $d$, sampling time steps $T$ and steps length $\mathrm{d}t$.
+Parameter: Gaussian process $\mathcal{GP}(0,K)$ for sampling $u_{0}$, noise level $\sigma$.
+Output: new airfoil $\{y_{i}\}$ at resolution $d$.
+1:  Inversely sample original airfoil function $\{u_{1,i}\}$ through $v_{\theta}$ and get its latent function $\{u_{0,i}\}$.
+2:  Set $\{a_{\theta,i}\}=\{u_{0,i}\}$.
+3:  while finetuning… do
+4:     Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
+[…]
+6:     Compute gradient and update $\theta$.
+7:  end while
+8:  Sample $\{u_{1,i}^{{}^{\prime}}\}$ through $v_{\theta}$ from $\{a_{\theta,i}\}$.
+9:  return $\{y_{i}\}=\{u_{1,i}^{{}^{\prime}}\}$
+Table 7: Training hyperparameters and model parameters in conditional airfoil generation tasks
 Training hyperparameters Value
-max learning rate 5e-6
-batch size 512
-optimizer used in pre-training and finetuning Adam
-train max iterations 1,000,000
-ODE Solver time steps 100
+Max learning rate $5\times 10^{-6}$
+Batch size $1024$
+Batch size (AF-200K) $2048$
+Optimizer (pre-training and fine-tuning) Adam
+Optimizer scheduler Cosine Annealing
+Warmup iterations $2000$
+Max training iterations $1000000$
+Max training iterations (AF-200K) $2000000$
+ODE solver time steps $10$
+ODE solver type Euler
 Model parameters Value
-Fourier neural operator layers 8
-Fourier neural operator modes 48
+Fourier neural operator layers $6$
+Fourier neural operator modes $64$
 Fourier neural operator hidden channels $256$
-Matérn kernel $\nu$ 2.5
-Matérn kernel $l$ 0.03
-We adopted the same conditions and symbols as those in AFBench.
-The parameters below, including the leading-edge radius, the curvature at the maximum thickness of the upper and lower surfaces, and the trailing-edge angle of the airfoil, were obtained through fitting.
-Note that our implementation of the angle calculation differs slightly from that of AFBench, as we found that the original fitting method was not invariant when the sampling precision increased.
-Appendix B Algorithm Details
-The algorithm for training airfoil generative models is shown in Algorithm 2, and inference using a trained model is shown in Algorithm 3.
-In the actual implementation, we utilized Optimal Transport techniques and employed the Matérn kernel as the kernel function for the Gaussian process in the latent space of the generative model.
+Gaussian process kernel Matérn
+Kernel noise scale $n$ $1$
+Kernel order $\nu$ $2.5$
+Kernel characteristic length $l$ $0.03$
+[…]
+Appendix B Algorithm Implementation Details
+The algorithm for training the airfoil generative model is presented in Algorithm 1, and the inference procedure using a trained model is described in Algorithm 2.
+In our implementation, we employed Optimal Flow Matching and utilized a Gaussian process with specific kernel functions, such as the RBF kernel or the Matérn kernel, in the latent space of the generative model.
 Appendix C Training Details
-In most of the experiments, we followed the hyperparameters listed in the table below to train the models, including both the pre-training and fine-tuning stages.
-On the AF-200K dataset, we increased the maximum number of iterations to 2,000,000 and the batch size to 2048.
-For the number of time steps used to solve the ODE, we tested reconstruction metrics on the super dataset with 10, 100, and 1000 steps, and ultimately chose 100 steps as the default setting for all other experiments, except for the fine-tuning stage, where we used 10 steps for training efficiency.
+In most of the experiments, we followed the hyperparameters listed in the Table 7 below to train the models, including both the pre-training and fine-tuning stages.
+On the AF-200K dataset, we increased the maximum number of iterations to $2000000$ and the batch size to $2048$.
+For the number of time steps used to solve the ODE, we use $10$ steps for relative balance between fast sampling speed and good performance, except for the fine-tuning stage, where we used 10 steps for training efficiency.
+Appendix D More Experiments
+We conduct a comprehensive ablation study on the choice of Gaussian process kernel types, model modes of Fourier Neural Operators, and characteristic lengths, as shown in Table 8 for the Matérn kernel, Table 9 for the radial basis function (RBF) kernel, and Table 10 for the white noise kernel, which is commonly used in diffusion or flow models in finite-dimensional spaces.
+In general, appropriate selections of these settings can significantly enhance model performance.
+Specifically, the Gaussian prior should exclude function spaces where the data is unlikely to reside, making the prior space as restricted as possible to facilitate effective model training.
+For example, a dataset with high smoothness is unlikely to belong to the function space generated by a white noise prior but is more plausibly sampled from a Gaussian prior with an RBF kernel.
+In Table 9, we observe that the RBF kernel exhibits a similar tendency to the Matérn kernel, with both achieving optimal performance at the characteristic length $0.03$.
+As for the white noise kernel in Table 10, since its characteristic length is effectively $0$, we studied the influence of the noise scale $n$ over the range $[1.0,2.0,4.0]$.
+We found that excessively large diffusion white noise negatively impacts model performance, likely because such noise scales project the function into a larger white noise space, making it more difficult for the neural operator to learn the intrinsic dynamics.
+Table 8: FuncGenFoil model performance with different FNO mode and Matérn kernel for conditional airfoil generation.
+[…]
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+$32$ $2.5$ $0.01$ 0.62 6.06 $\mathbf{0.09}$ 200.4 3.58 0.09 185.3 $\mathbf{0.07}$ 0.12 26.7 29.1 41.1 2.51 -105.6 1.02
+$0.02$ 0.52 5.10 0.10 97.4 3.26 0.11 107.0 0.08 0.12 26.2 $\mathbf{27.6}$ 24.3 2.22 -105.7 $\mathbf{0.99}$
+$0.03$ 0.54 $\mathbf{4.78}$ 0.11 $\mathbf{82.3}$ 3.00 0.11 93.2 0.09 0.12 $\mathbf{26.0}$ 27.7 $\mathbf{21.6}$ $\mathbf{2.18}$105.6 $\mathbf{0.99}$
+$3.5$ $0.01$ 0.63 5.64 $\mathbf{0.09}$ 158.3 3.46 $\mathbf{0.08}$ 156.9 $\mathbf{0.07}$ 0.12 27.6 29.2 34.7 2.41 -105.8 1.01
+$0.02$ $\mathbf{0.48}$ 5.08 0.10 89.4 3.12 0.10 97.5 0.08 0.12 26.8 28.2 22.8 $\mathbf{2.18}$105.9 $\mathbf{0.99}$
+$0.03$ 0.53 5.17 0.11 83.5 $\mathbf{2.90}$ 0.11 $\mathbf{89.5}$ 0.10 0.14 29.4 30.7 22.0 2.28 -105.2 $\mathbf{0.99}$
+$64$ $1.5$ $0.01$ 0.82 8.57 0.10 370.2 5.21 0.09 342.5 0.08 0.12 36.3 37.5 72.9 3.30 -104.0 1.09
+$0.02$ 0.65 8.26 0.13 222.9 4.97 0.11 219.6 0.08 0.13 37.4 35.9 48.2 3.12 -103.6 1.04
+$0.03$ 0.71 8.18 0.14 183.7 4.82 0.12 207.2 0.09 0.14 34.6 36.6 43.3 3.13 -104.0 1.02
+$2.5$ $0.01$ 0.80 8.77 0.10 291.2 5.23 0.10 277.6 0.08 0.12 35.8 37.7 59.8 3.20 -103.7 1.06
+$0.02$ 0.68 8.66 0.13 192.4 4.78 0.11 182.8 0.09 0.12 36.2 37.8 42.2 3.06 -103.5 1.02
+$0.03$ 0.71 8.23 0.13 201.3 4.72 0.12 174.2 0.09 0.14 34.2 36.7 41.9 3.08 -103.9 1.01
+$3.5$ $0.01$ 0.88 8.80 0.11 269.3 5.26 0.09 248.9 0.08 0.12 36.5 38.0 55.3 3.20 -103.4 1.05
+$0.02$ 0.72 8.52 0.11 177.1 4.84 0.11 174.6 0.08 0.13 33.2 37.4 39.7 2.98 -103.6 1.02
+$0.03$ 0.66 7.80 0.12 161.1 4.54 0.11 164.2 0.09 0.15 35.1 38.3 37.5 2.96 -103.4 1.01
+$0.04$ 0.82 8.19 0.13 170.0 4.45 0.13 159.7 0.09 0.16 36.6 34.8 37.7 3.11 -103.6 1.01
+$0.06$ 0.87 8.82 0.16 183.0 5.01 0.13 183.9 0.10 0.17 41.8 41.2 42.3 3.44 -102.8 1.02
+$0.12$ 1.39 10.7 0.20 261.0 6.23 0.15 230.3 0.12 0.19 53.4 44.9 55.3 4.27 $\mathbf{-99.9}$ 1.05
+$128$ $2.5$ $0.01$ 1.09 10.4 0.11 708.5 7.13 0.10 613.5 0.09 $\mathbf{0.11}$ 40.7 41.1 129.3 4.16 -102.8 1.27
+$0.02$ 1.15 10.0 0.11 577.5 6.83 0.11 570.2 0.09 0.14 37.8 42.3 113.3 4.14 -103.4 1.20
+$0.03$ 1.05 10.1 0.12 598.0 6.48 0.14 487.7 0.09 0.15 36.9 40.1 107.3 4.22 -103.0 1.23
+$3.5$ $0.01$ 1.04 10.3 0.10 628.9 7.05 0.09 576.8 0.08 0.12 39.3 42.0 118.7 4.05 -102.9 1.23
+$0.02$ 0.93 9.94 0.12 560.6 6.81 0.10 484.6 0.09 0.15 41.5 40.8 104.1 4.05 -103.2 1.19
+$0.03$ 1.11 9.93 0.11 548.2 6.49 0.12 520.3 0.09 0.16 38.6 39.0 105.8 4.16 -103.1 1.23
+Table 9: FuncGenFoil model performance with different RBF kernel for conditional airfoil generation.
+[…]
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+$64$ $0.001$ 1.26 $\mathbf{7.60}$ $\mathbf{0.09}$ 1224 5.34 $\mathbf{0.07}$ 1046 $\mathbf{0.06}$ $\mathbf{0.08}$ 24.9 $\mathbf{28.6}$ 213 3.68 -103.7 1.42
+$0.01$ 0.77 8.67 0.12 223.6 5.05 0.09 174.7 0.08 0.12 35.2 37.1 47.0 3.06 -103.6 1.03
+$0.02$ $\mathbf{0.62}$ 8.33 0.12 162.2 4.59 0.11 165.9 0.09 0.17 38.0 42.5 38.4 3.04 -103.2 $\mathbf{1.01}$
+$0.03$ 0.66 7.80 0.12 $\mathbf{147.9}$ $\mathbf{4.43}$ 0.12 $\mathbf{154.2}$ 0.09 0.15 $\mathbf{33.2}$ 37.9 $\mathbf{35.1}$ $\mathbf{2.93}$103.6 1.00
+$0.04$ 0.84 8.79 0.13 197.1 4.59 0.12 168.4 0.09 0.17 35.8 37.9 41.3 3.20 -103.2 1.02
+$0.06$ 1.15 9.97 0.15 220.3 6.21 0.14 215.5 0.10 0.16 38.3 44.2 48.7 3.73 -99.9 1.05
+$0.12$ 1.12 10.9 0.17 230.5 7.71 0.17 265.6 0.11 0.18 40.0 42.3 54.4 4.14 $\mathbf{-95.1}$ 1.08
+Table 10: FuncGenFoil model performance with white noise kernel of different noise scale $n$ for conditional airfoil generation.
+[…]
+$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma_{9}$ $\sigma_{10}$ $\sigma_{11}$ $\bar{\sigma}_{a}$ $\bar{\sigma}_{g}$
+$64$ 1 1.15 7.18 0.09 913.8 5.16 0.07 964.5 0.06 0.08 26.2 26.1 176.8 3.44 -104.7 1.39
+2 1.46 7.98 0.10 1274 5.94 0.08 1268 0.08 0.10 30.6 32.4 238.3 4.17 -102.7 1.64
+4 2.27 9.14 0.12 2015 7.24 0.11 2059 0.09 0.14 41.5 41.4 379.7 5.61 -99.9 2.09
+Appendix E Limitations
+Although we have taken a step toward general object shape modeling in function space, the focus of this paper is primarily on airfoils or aircraft wings, which have relatively simple geometries and smooth surfaces yet are highly significant for aerodynamic performance.
+Thus, our current approach has limitations regarding the scope of geometry.
+In future studies, if we aim to extend our method to modeling entire aircraft or dealing with objects of general shapes, substantial theoretical work and experimental analysis will be required, especially when a suitable coordinate system is unavailable for describing complex, irregular, or non-smooth geometries.
+This direction is part of our ongoing efforts.
+Appendix F Societal Impacts
+This work will accelerate research in AI for Science and Engineering and generally has positive social impacts.
+On the other hand, AI models for engineering design and optimization could accelerate technological diffusion within society, potentially raising issues related to intellectual property or unauthorized technology transfer to entities or individuals intending to design specific engineering products or construct items harmful to society.